### v1

# Improving Matrix Exponential for Generative AI Flows: A Taylor-Based Approach Beyond Paterson–Stockmeyer

Abstract
The matrix exponential is a fundamental operator in scientific computing and system simulation, with applications ranging from control theory and quantum mechanics to modern generative machine learning. While Padé approximants combined with scaling and squaring have long served as the standard, recent Taylor-based methods, which utilize polynomial evaluation schemes that surpass the classical Paterson–Stockmeyer technique, offer superior accuracy and reduced computational complexity. This paper presents an optimized Taylor-based algorithm for the matrix exponential, specifically designed for the high-throughput requirements of generative AI flows. We provide a rigorous error analysis and develop a dynamic selection strategy for the Taylor order and scaling factor to minimize computational effort under a prescribed error tolerance. Extensive numerical experiments demonstrate that our approach provides significant acceleration and maintains high numerical stability compared to existing state-of-the-art implementations. These results establish the proposed method as a highly efficient tool for large-scale generative modeling.
keywords:
Matrix exponential , scaling and squaring , Taylor approximation , Padé approximant , efficient matrix polynomial evaluation , Algorithmic Efficiency , Generative AI flows.
2020 MSC:
65F60 , 41A10 , 65D15 , 41-04 , 68T07 , 65Y20
journal: Elsevier
1 Introduction
The computation of matrix functions, particularly the matrix exponential, has been extensively studied due to its wide range of applications in science and engineering [1, 2]. Among the various numerical methods developed, Padé approximants have traditionally been favored over Taylor series due to their superior efficiency for comparable accuracy [23]. However, despite its widespread adoption, the standard Padé-based approach often incurs unnecessary computational overhead. Recent work, such as [26], has demonstrated that truncated Taylor series can represent a simpler and more efficient alternative, frequently yielding substantial speedups over the Padé-based scaling and squaring implementations found in standard software routines like MATLAB’s expm.
In fact, recent advances have shown that Taylor-based methods can outperform Padé approximants in both accuracy and efficiency. The Taylor algorithm introduced in [3] improved upon classical Padé-based methods [23], while subsequent research [4, 5, 6] refined the associated error analysis and reduced computational costs.
A significant breakthrough in this area was the development of advanced matrix polynomial evaluation techniques [9] that surpass the classical Paterson–Stockmeyer method [13]. These techniques, predicated on structured multiplications and linear combinations of matrix polynomials, enable higher-order Taylor approximations with reduced computational effort. For instance, the method presented in [22] introduced $m+$ approximations, where the first $m$ terms match the Taylor series exactly, while additional terms are optimized to enhance accuracy. These schemes, such as the $m=15+$ case introduced in [9, Ex. 5.1], achieve greater precision than standard Taylor polynomials of equivalent computational cost. Furthermore, recent optimizations for determining evaluation formula coefficients [18] have shown promising results for broader matrix function computations [18, 19].
The cost of these polynomial approximation methods is dominated by matrix multiplications, denoted by $M$. In practice, these operations scale as $\mathcal{O}(n^{3})$ for $n\times n$ matrices, whereas matrix additions and scalar multiplications are of order $\mathcal{O}(n^{2})$. In rational methods, matrix inverses are computed by solving linear systems, the cost of which can also be expressed in terms of matrix products as [12, App. C]:
$D\approx\frac{4}{3}M.$ (1)
Consequently, throughout this paper, the cost of both rational and polynomial approximations is quantified in terms of $M$.
Table 1 compares the cost and achievable approximation orders of various state-of-the-art methods, including Padé approximants, Paterson–Stockmeyer evaluation, and mixed rational-polynomial schemes [11, 22, 14]. The results highlight the superior efficiency of modern polynomial evaluation strategies, which facilitate higher-order approximations using fewer matrix multiplications.
In this paper, we propose an optimized implementation of the matrix exponential based on the evaluation methods described in [9]. Our approach is specifically tailored for the high-throughput requirements of Generative AI flows, where the repeated evaluation of the matrix exponential often constitutes a primary computational bottleneck. Unlike traditional software routines optimized for fixed IEEE precision using pre-computed thresholds [21, 14, 22, 15], or flexible implementations that depend on external specialized libraries [16], we introduce a self-contained framework that dynamically selects the Taylor order and scaling factor for any user-defined tolerance $\varepsilon$. While previous dynamic methods have utilized classical Paterson–Stockmeyer schemes with more complex scaling strategies [17], the proposed algorithm leverages higher-order polynomial evaluations with a simplified operational logic. This provides a portable, library-independent solution that balances numerical stability with significant reductions in execution time for large-scale generative models.
Polynomial evaluation cost $3M$ $4M$ $5M$ $6M$ $7M$
Approx. order $m$ Paterson–Stockmeyer [13] 6 9 12 16 20
Approx. order $m$ [14] 8 12 18 22
Approx. order $m$ [22] 8 $15+$ $21+$ 24 30
Mixed rational polynomial approx. cost $3.33M$ $4.33M$ $5.33M$ $6M$ $7M$
Approx. order from method [11, Tab. 3] 9 12 16 21 28
Padé evaluation cost $3.33M$ $4.33M$ $5.33M$ $6.33M$ $7.33M$
Approx. order Padé method [23, Tab. 2.2] 6 10 14 18 26
Table 1: Cost and available order of approximation with the polynomial evaluation methods from Paterson–Stockmeyer [13], Sastre–Ibáñez–Defez [22, Sec. 3], Bader–Blanes–Casas [14, Sec. 4], the mixed rational and polynomial approximation method from Sastre [11, Tab. 3] and the Padé approximation method from Higham [23, Tab. 2.2].
Throughout this manuscript, $\mathbb{C}^{n\times n}$ denotes the set of complex matrices of size $n\times n$, $I$ denotes the identity matrix for this set, $\rho(A)$ is the spectral radius of matrix $A$, and $\mathbb{N}$ denotes the set of positive integers. The matrix norm $\left\|\cdot\right\|$ denotes any subordinate matrix norm; in particular $\left\|\cdot\right\|_{1}$ is the 1-norm. The symbols $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ denote the smallest following and the largest previous integer, respectively.
This paper begins in Section 2 by reviewing the role and challenges associated with the matrix exponential in generative flow models. Section 3 introduces the proposed computational methodology, which includes new evaluation formulas for the Taylor matrix polynomial approximation (Subsection 3.1) based on [9, 10], a discussion of the scaling and squaring error analysis (Subsection 3.2), and the complete algorithmic procedure (Subsection 3.3). Subsequently, Sections 4 and 5 provide the numerical results that assess the algorithm’s performance, and Section 6 presents the conclusions.
2 The matrix exponential for generative flows
2.1 Generative models
In recent years, generative models have become very popular. They aim to use training data to learn the probability distribution of a dataset. The resulting model of the distribution can be sampled to obtain new data points according to their estimated probability. This has proven useful in a variety of tasks, including music generation [27, 28], denoising [29], and representation learning [30].
There are a variety of approaches to train these probabilistic models. Generative adversarial networks (GANs) [31] train two models with adversarial objectives. Variational Autoencoders (VAEs) [32] take a bayesian approach to modelling the data with a prior. Denoising diffusion models [33] aim to model a denoising process (sometimes called score matching) to progressively transform a noise signal into a data point. While these techniques have all been applied succesfully to generative tasks, they unfortunately lack invertibility. Flow generative models were developed to have an invertible mapping from a noise distribution to a data distribution.
2.2 Generative flows
For random variable $x\in\mathbb{R}^{n}$ representing data with probability distribution $p(x)$ and random variable $z\in\mathbb{R}^{n}$ with a known probability density function $p(z)$, generative flows aim to model an invertible transformation of $x$ into $z=f(x)$. By the change of variable formula, the relationship between $p(x)$ and $p(z)$ is:
$\log p(x)=\log p(z)+\log\left|\frac{\delta f(x)}{\delta x}\right|,$ (2)
where $\frac{\delta f(x)}{\delta x}$ is the Jacobian of $f$. Assume that there are $k$ number of invertible functions. Then, the log-likelihood of $x$ is given by
$\log p(x)=\log p(z)+\sum^{k}_{i=1}\log\left|\frac{\delta h_{i}}{\delta h_{i-1}}\right|,$ (3)
where $h_{i}=f_{i}(h_{i-1})$.
In generative flow models [25], the transformation function $f$ is typically implemented as a composition of invertible functions, often parameterized as $f=W_{m}\,\phi(W_{m-1}\,\phi(\cdots\phi(W_{1}x)))$, where $W_{i}$ are weight matrices and $\phi$ is a strictly monotonic activation function. Ensuring invertibility of each $W_{i}$ during training is challenging, and Calculating the determinant of a general matrix is computationally intensive, with a complexity of $O(n^{3})$ for an $n\times n$ matrix. A common workaround is to constrain $W_{i}$ to be triangular, which simplifies determinant computation but limits expressiveness and parallelizability.
Over the years, there have been multiple approaches to modeling normalizing flows that address this computationally expensive log determinant problem. One solution is to model $f$ autoregresively [40], calculating the determinant in $\mathcal{O}(n)$ time. Contractive residual layers have been used to implement invertible transforms [42]. They result in unbiased density estimators at the cost of long computation time. Multiple approaches use coupling layers [39, 41] to implement unbiased reversible transformations with a triangular matrix for a traceable determinant calculation.
One approach [25] proposed replacing each weight matrix $W_{i}$ with its matrix exponential $e^{W_{i}}$. This guarantees invertibility, improves expressiveness over triangular matrices, and enables efficient computation of both the inverse and the log-determinant. Specifically, the inverse is given by $(e^{W})^{-1}=e^{-W}$, and the log-determinant simplifies to $\log\det(e^{W})=\operatorname{Tr}(W)$, reducing the computational cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
The matrix exponential maps $\mathbb{M}_{n}(\mathbb{R})$ to $\mathrm{GL}_{n}(\mathbb{R})^{+}$, the group of invertible matrices with positive determinant. Although not surjective, its image is rich enough for practical purposes, and any matrix in $\mathrm{GL}_{n}(\mathbb{R})^{+}$ can be expressed as a product of matrix exponentials. This makes the matrix exponential a powerful tool for constructing expressive and invertible neural architectures in generative modeling [25].
Algorithm 1 Algorithm for Computing Matrix Exponential
1:Weight matrix $W$, error tolerance $\varepsilon$
2:Matrix exponential $e^{W}$
3:Choose the smallest non-negative integer $s$ such that $\|W\|_{1}/(2^{s})<\frac{1}{2}$
[…]
5:$X=I$
6:$Y=W$
7:$k=2$
8:while $\|Y\|_{1}>\varepsilon$ do
9:  $X=X+Y$
10:  $Y=W\cdot Y/k$
11:  $k=k+1$
12:end while
13:for $i=1$ to $s$ do
14:  $Y=Y\cdot Y$
15:end for
16:return $Y$
For the sake of completeness the algorithm for the computation of the matrix exponential described in [25, Sec. 3.2] is reproduced here as Algorithm 1. It relies on the matrix exponential Taylor series expansion [2, Sec. 3], denoted by $T_{m}(W)$ and given by
$T_{m}(W)=\sum_{i=0}^{m}\frac{W^{i}}{i!},$ (4)
where terms are successively summed until the norm of the last term falls below a prescribed tolerance $\varepsilon$. This approach uses a forward absolute error criterion based on an error bound for the Taylor series remainder for approximating the matrix exponential, denoted by $R_{m}(W)$ and given by
$R_{m}(W)=e^{W}-T_{m}(W)=\sum_{i=m+1}^{\infty}\frac{W^{i}}{i!},$ (5)
which is bounded by [34]:
$\|R_{m}(W)\|_{1}\leq\frac{\|W\|_{1}^{m+1}}{(m+1)!}\left(\frac{1}{1-\frac{\|W\|_{1}}{m+2}}\right).$ (6)
Note that the truncated Taylor series is particularly well-suited for scenarios with limited computational resources, restricted access to external libraries or large-scale problems involving well-behaved matrices [26, p. 52]. According to [25, Sec. 3.2], the weight matrices in neural networks, denoted by $W$, are often small in norm such that $|W|_{1}$ is also small. This property makes the incorporation of the matrix exponential into neural network architectures computationally feasible. Exploiting the identity $e^{W}=(e^{W/2^{s}})^{2^{s}}$, in Algorithm 1 the matrix is first scaled to ensure that $|W/2^{s}|<1/2$, after which the exponential is computed and subsequently squared $2^{s}$ times. This scaling and squaring strategy further reduces the computational cost while maintaining numerical stability; see [25, Alg. 1]. The computational cost of evaluating a Taylor approximation of degree $m$ using [25, Alg. 1], denoted by $T_{m}(A)$, is measured in terms of matrix multiplications. Let $C_{\text{orig}}$ represent the cost of evaluating the unscaled Taylor polynomial in Algorithm [25, Alg. 1], which is given by
$C_{\text{orig}}=m-1.$ (7)
When scaling is applied, the total cost becomes (s+m1)M(s+m-1)M, where $s$ is the scaling parameter. As reported in [25, Sec. 6.4], a tolerance of $\varepsilon=10^{-8}$ is used. Under this setting, the total number of matrix products, given by s+m1s+m-1, does not exceed 11, with an average value of 9.28 observed during training [25, Tab. 6].
Cost (7) makes Algorithm 1 unsuitable for high-dimensional data [25, Sec. 3.2]. To address this limitation, a second method is proposed that combines matrix exponentials with neural networks, using a low-rank parameterization of the weight matrix $W=A_{1}A_{2}$, where $A_{1}\in\mathbb{R}^{n\times t}$ and $A_{2}\in\mathbb{R}^{t\times n}$.
By defining $V=A_{2}A_{1}$, and using the associative property of matrix multiplication, the matrix exponential $e^{W}$ can be approximated via a truncated series expansion of $V$:
$e^{W}\approx I+A_{1}\left(\sum_{i=0}^{m}\frac{V^{i}}{(i+1)!}\right)A_{2},$ (8)
where the truncation error bound of the matrix series of $V$, denoted here by $R^{\prime}_{m}(V)$, is given by (8) from [25, Sec. 3.2] reproduced here
$\left\|R^{\prime}_{m}(V)\right\|_{1}=\left\|\sum_{i=m+1}^{\infty}\frac{V^{i}}{(i+1)!}\right\|_{1}\leq\left(\frac{\|V\|_{1}^{m+1}}{(m+2)!}\right)\left(\frac{1}{1-\frac{\|V\|_{1}}{m+3}}\right).$ (9)
This reduces the computational cost to $\mathcal{O}(t^{3})$ since $V\in\mathbb{R}^{t\times t}$. The rank of $W$ is at most $t$, so choosing a smaller $t$ improves efficiency but may reduce expressiveness. Thus, there is a trade-off between computational cost and model capacity.
The approximation method is similar to Algorithm 1, with modifications: set the scale coefficient $s:=0$, change line 4 to $Y:=W/2$, and line 5 to $k:=3$ [25, Sec. 3.2].
While the cost of the algorithms proposed in [25, Sec. 3.2] represent a reasonable computational cost, further reductions are possible by optimizing the polynomial evaluation strategy.
3 Proposed method
Our implementation is specifically tailored for this generative flow setting. It builds upon efficient polynomial evaluation techniques, extending beyond the classical Paterson–Stockmeyer method as introduced in [9], to substantially reduce the number of matrix products required for a specified approximation order. Furthermore, the algorithm incorporates a dynamic procedure that optimally selects the Taylor series degree ($m$) and the scaling factor ($s$) to minimize the overall computational cost. This selection ensures that the error tolerance is strictly maintained, chosen to be no less than the unit roundoff ($u$) of the arithmetic used, while also strategically reducing the risk of overscaling [21, Sec. 1].
Numerical experiments confirm that our approach consistently achieves lower average computational costs than the baseline method in [25], particularly in the context of generative AI flows, where evaluations of thousands of matrix exponentials are required. This improvement translates into faster training times and reduced resource consumption, making our method suitable for large-scale machine learning applications.
3.1 Evaluation formulas of the Taylor-based matrix exponential approximations
This work adopts the evaluation formulas from [22] to compute Taylor-based approximations for orders up to $m=15+$. Table 1 demonstrates that these schemes significantly outperform the Paterson–Stockmeyer method; for a fixed computational budget, they achieve orders $8$ and $15+$, whereas the classical approach only reaches orders $6$ and $9$, respectively. Moreover, these schemes represent the most efficient approximations identified in Table 1, outperforming both the state-of-the-art Taylor [14] and Padé [23] methods. Although [22] also provides formulas for higher-order Taylor approximations, given the low error tolerance required in our application, these higher-order variants do not offer a reduction in computational cost and are therefore excluded from consideration. For completeness, the formulas corresponding to the selected approximation orders are reproduced below.
$T_{1}(A)$ $=$ $A+I,$ (10)
$T_{2}(A)$ $=$ $A^{2}/2+A+I,$ (11)
$T_{4}(A)$ $=$ $((A^{2}/4+A)/3+I)A^{2}/2+A+I,$ (12)
$y_{02}(A)$ $=$ $A^{2}(c_{1}A^{2}+c_{2}A),$ (13)
$T_{8}(A)$ $=$ $(y_{02}(A)+c_{3}A^{2}+c_{4}A)(y_{02}(A)+c_{5}A^{2})$
$+c_{6}y_{02}(A)+A^{2}/2+A+I,$
where the coefficients $c_{i}$, for $i=1,\,2\,,\ldots,\,6$, are listed in Table 2, rounded to IEEE double-precision floating-point arithmetic.
[…]
Table 2: Coefficients for the Taylor approximation of the matrix exponential of order $m=8$, computed using formulas (13) and (3.1)
$y_{02}(A)$ $=$ $A^{2}(c_{1}A^{2}+c_{2}A),$ (15)
$y_{12}(A)$ $=$ $(y_{02}(A)+c_{3}A^{2}+c_{4}A)(y_{02}(A)+c_{5}A^{2})+c_{6}y_{02}(A)+c_{7}A^{2},$ (16)
$y_{22}(A)$ $=$ $(y_{12}(A)+c_{8}A^{2}+c_{9}A)(y_{12}(A)+c_{10}y_{02}(A)+c_{11}A)$ (17)
$+c_{12}y_{12}(A)+c_{13}y_{02}(A)+c_{14}A^{2}+c_{15}A+c_{16}I,$
where the coefficients $c_{i}$, for $i=1,\,2,\,\ldots,\,16$, are listed in Table 3, also rounded to IEEE double-precision floating-point arithmetic.
[…]
Table 3: Coefficients of $y_{02}$, $y_{12}$, and $y_{22}$ from equations (15)-(17) used in the Taylor-based approximation of the matrix exponential of order $m=15+$.
Note that
  1. 1.
    $T_{1}(A)$ and $T_{2}(A)$ in (10) correspond to the direct Taylor evaluation formulas of orders 1 and 2, respectively, while $T_{4}(A)$ corresponds to the Paterson–Stockmeyer evaluation scheme [13] for order $m=4$.
  2. 2.
    The cost of evaluating $T_{8}(A)$ using formulas (13) and (3.1) is $3M$: one multiplication for computing $A^{2}$, one for evaluating (13), and one for evaluating (3.1). In contrast, computing the Taylor approximation of degree $m=8$ using the original algorithm from [25, Alg. 1], as given by (7), requires $7M$, which is more than twice as expensive.
  3. 3.
    Recall that, in exact arithmetic, equations (15) and (16) from [22] show that $y_{22}(A)$ in (17) satisfies
    $y_{22}(A)=T_{15}(A)+b_{16}A^{16},$ (18)
    and, therefore, the approximation remainder is given by
    $R_{15+}(W)=e^{W}-y_{22}(W)=\left(\frac{1}{16!}-b_{16}\right)W^{16}+\sum_{i=17}^{\infty}\frac{W^{i}}{i!},$ (19)
    where, from (15)–(17), it can be shown that the coefficient is $b_{16}=c_{1}^{4}$. Using the value of $c_{1}$ from Table 3, we obtain
    $b_{16}=c_{1}^{4}\approx 2.608368698098256\times 10^{-14}.$ (20)
    This value differs from the exact Taylor coefficient $1/16!$ by a relative error of approximately $|b_{16}-1/16!|\cdot 16!\approx 0.454$. Consequently, $y_{22}(A)$ provides a more accurate approximation than $T_{15}(A)$ in exact arithmetic, although it remains less accurate than the full Taylor approximation $T_{16}(A)$ as demonstrated in [22]. Approximations of this type are referred to in [22] as $T_{15+}(A)$. Similarly, in this case, the cost of evaluating $y_{22}(A)$ using formulas (15)-(17) is $4M$.
3.2 Error analysis
This section details the self-contained framework for the dynamic selection of the Taylor order $m$ and the scaling parameter $s$. Unlike fixed-precision algorithms that rely on pre-computed thresholds for specific IEEE formats or error tolerances [21, 14, 22, 15], the proposed approach derives these parameters directly for a user-specified tolerance $\varepsilon$. By integrating advanced polynomial evaluation formulas into the error bound analysis, we achieve a more streamlined selection logic than previous dynamic methods [17].
In particular, we provide a refined version of the absolute error analysis from (6) specifically for nonnormal matrices. For such matrices, the relation (1.5) from [21] states that
$\rho(A)\leq\|A^{k}\|^{1/k}\leq\|A\|,\quad k=1,2,\dots,\infty.$ (21)
Consequently, the inequality
$\|A^{k}\|\leq\|A\|^{k},\quad k=1,2,\dots,\infty,$ (22)
can be significantly strict, i.e., $\|A^{k}\|\ll\|A\|^{k}$, as discussed in [21, Sec. 1]. This property is exploited in [5, Th. 1.1] to derive tighter bounds for the matrix power series, ensuring efficiency without the overhead of overscaling. For completeness of the exposition, the theorem is reproduced below:
Theorem 1.
[…]
$\alpha_{p}=\max\{a_{k}^{\frac{1}{k}}:k=p,l,l+1,l+2,\ldots,p_{0}-1,p_{0}+1,p_{0}+2,\ldots,l+p-1\},$ (23)
then
$h_{l}(A)||\leq\tilde{h}_{l}(\alpha_{p}).$ (24)
Using Theorem 1 the following theorem provides a sharper bound than (6) given originally by [34] for the remainder $R_{m}(W)$ of the matrix exponential Taylor series of nonnormal matrices.
Theorem 2.
Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. Suppose that for each $k\in\mathbb{N}$, there exists an upper bound $a_{k}$ such that $\|A^{k}\|_{1}\leq a_{k}$. Let $p\in\mathbb{N}$ with $1\leq p\leq m+1$, and let $p_{0}$ be the multiple of $p$ satisfying $m+1\leq p_{0}\leq m+1+p$. Define
$\alpha_{p}=\max\left\{a_{k}^{1/k}:k=p,m+1,m+2,\ldots,p_{0}-1,p_{0}+1,p_{0}+2,\ldots,m+1+p\right\}.$ (25)
Then the remainder of the matrix exponential Taylor approximation of degree $m$,
$R_{m}(A)=e^{A}-\sum_{k=0}^{m}\frac{A^{k}}{k!}=\sum_{k=m+1}^{\infty}\frac{A^{k}}{k!},$ (26)
satisfies the bound
$\|R_{m}(A)\|_{1}<\frac{\alpha_{p}^{m+1}}{(m+1)!}\cdot\left(\frac{1}{1-\frac{\alpha_{p}}{m+2}}\right),$ (27)
provided that condition
$\alpha_{p}<m+2,$ (28)
is fulfilled.
Proof.
Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. The remainder of the matrix exponential Taylor approximation of degree $m$ given by (5)
$R_{m}(A)=e^{A}-T_{m}(A)=\sum_{k=m+1}^{\infty}\frac{A^{k}}{k!}.$
Define the scalar power series
$h_{m+1}(x)=\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)=\sum_{k=m+1}^{\infty}\left|\frac{1}{k!}\right|x^{k},$
which has infinite radius of convergence. Suppose that for each $k\in\mathbb{N}$, there exists an upper bound $a_{k}\geq 0$ such that $\|A^{k}\|_{1}\leq a_{k}$.
Let $p\in\mathbb{N}$ with $1\leq p\leq m+1$, and let $p_{0}$ be the multiple of $p$ satisfying $m+1\leq p_{0}\leq m+1+p$. Define
$\alpha_{p}=\max\left\{a_{k}^{1/k}:k=p,m+1,m+2,\ldots,p_{0}-1,p_{0}+1,p_{0}+2,\ldots,m+1+p\right\}.$
By Theorem 1, we have
$\|R_{m}(A)\|_{1}=\left\|\sum_{k=m+1}^{\infty}\frac{A^{k}}{k!}\right\|_{1}\leq\sum_{k=m+1}^{\infty}\frac{\|A^{k}\|_{1}}{k!}\leq\sum_{k=m+1}^{\infty}\frac{a_{k}}{k!}\leq\tilde{h}_{m+1}(\alpha_{p}).$ (29)
To bound $\tilde{h}_{m+1}(\alpha_{p})$, note that factoring out the first term one gets
$\tilde{h}_{m+1}(x)=\frac{x^{m+1}}{(m+1)!}\cdot\sum_{k=0}^{\infty}\frac{x^{k}}{(m+2)(m+3)\cdots(m+1+k)}.$
Note that for each $k\geq 0$,
$(m+2)(m+3)\cdots(m+1+k)>(m+2)^{k},$
so
$\frac{x^{k}}{(m+2)(m+3)\cdots(m+1+k)}<\left(\frac{x}{m+2}\right)^{k}.$
Therefore,
$\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\sum_{k=0}^{\infty}\left(\frac{x}{m+2}\right)^{k}.$
This is a geometric series with ratio $\frac{x}{m+2}$, and if $\frac{x}{m+2}<1$ it converges to
$\sum_{k=0}^{\infty}\left(\frac{x}{m+2}\right)^{k}=\frac{1}{1-\frac{x}{m+2}}.$
Hence,
$\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\frac{1}{1-\frac{x}{m+2}},\quad\text{for }x<m+2.$ (30)
Applying (29) and (30) with $x=\alpha_{p}<m+2$ then (27) holds. ∎
In the following, The matrix exponential scaling parameter, $s$, is determined based on the result of Theorem 2 to ensure that the Taylor series remainder remains below a predefined error tolerance. Let $W$ be the weight matrix for generative flows from Algorithm 1. Let $a_{k}$ be an upper bound for $\|W^{k}\|_{1}$ such that $\|W^{k}\|_{1}\leq a_{k}$. Let $p\in\mathbb{N}$ with $1\leq p\leq m+1$, and let $p_{0}$ be the multiple of $p$ satisfying $m+1\leq p_{0}\leq m+1+p$. Let $\alpha_{p}$ be defined by (25). Let’s scale the the matrix $W$ by the scaling parameter $s$ so that
$\frac{\|(W/2^{s})^{m+1}\|_{1}}{(m+1)!}\leq\frac{(\alpha_{p}/2^{s})^{m+1}}{(m+1)!}\leq\varepsilon,$ (31)
where $\varepsilon$ is the selected user error tolerance from Algorithm 1. Note that $\varepsilon$ must fulfill
$\varepsilon\geq u,$ (32)
where $u$ is the machine unit roundoff error in the working precision arithmetic, since errors below the machine roundoff error can’t be guaranteed in the working precision arithmetic computations. Then, if $s$ fulfills (31) one gets
$s\geq\frac{1}{m+1}\log_{2}\left(\frac{\alpha_{p}^{m+1}}{(m+1)!\cdot\varepsilon}\right),$ (33)
and then it follows that
$s=\left\lceil\frac{1}{m+1}\log_{2}\left(\frac{\alpha_{p}^{m+1}}{(m+1)!\cdot\varepsilon}\right)\right\rceil.$ (34)
Since $s$ is the lowest integer satisfying (34), then scaling the matrix $W$ by $s$ given by (34) it follows that
$\frac{\|(W/2^{s})^{m+1}\|_{1}}{(m+1)!}\leq\frac{(\alpha_{p}/2^{s})^{m+1}}{(m+1)!}=\varepsilon^{\prime}\leq\varepsilon,$ (35)
and by (27) from Theorem 2, if $\alpha_{p}/2^{s}<m+2$ one gets
$\|R_{m}(W/2^{s})\|_{1}\leq\varepsilon^{\prime}\frac{1}{1-\frac{\alpha_{p}/2^{s}}{m+2}},$ (36)
where $\alpha_{p}/2^{s}=\varepsilon^{\prime\frac{1}{{m+1}}}\leq\varepsilon^{\frac{1}{{m+1}}}$ and then
$\|R_{m}(W/2^{s})\|_{1}\leq\varepsilon\frac{1}{1-\frac{\varepsilon^{1/{m+1}}}{m+2}}.$ (37)
We verify the fulfillment of condition (28) by utilizing the error tolerance $\varepsilon=10^{-8}$ reported in [25, Sec. 6.4] and the selected Taylor degrees $m\in\{1,2,4,8,15\}$ (as defined in Section 3.1). For these parameters, it is easy to show that the condition is satisfied since the maximum value of the bound, $\alpha_{p}/2^{s}$, meets the requirement
$\alpha_{p}/2^{s}\leq\varepsilon^{\frac{1}{m+1}}=(10^{-8})^{\frac{1}{m+1}}<m+2,$
for each value of $m$. Subsequently, analyzing the maximum total error bound given by (36), we find it equals
$\varepsilon+1.75682\times 10^{-18},$
where the second term is approximately ten orders of magnitude smaller than the target tolerance $\varepsilon$. This effectively means the total error bound is dominated by $\varepsilon$, validating the scaling parameter selection derived from equation (34) under practical floating-point arithmetic.
To ensure that the bound is accomplished properly the proposed algorithm will compute the scaling parameter $s$ by using the term $m+1$ as stated here, and also the term $m+2$, selecting the largest scaling parameter among both of them. For the case $m=15+$, where the term $m+1=16$ is not exactly $1/16!$, selecting the maximum scaling parameter obtained with $m+1$ and also the term $m+2=17$, for which the exponential remainder term corresponds effectively to $1/17!$, will ensure a proper scaling parameter value.
The error analysis for (8) is analogous to the error analysis described here, and the following theorem can be proved similarly:
Theorem 3.
Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. Suppose that for each $k\in\mathbb{N}$, there exists an upper bound $a_{k}$ such that $\|A^{k}\|_{1}\leq a_{k}$. Let $p\in\mathbb{N}$ with $1\leq p\leq m+1$, and let $p_{0}$ be the multiple of $p$ satisfying $m+1\leq p_{0}\leq m+1+p$. Define
$\alpha_{p}=\max\left\{a_{k}^{1/k}:k=p,m+1,m+2,\ldots,p_{0}-1,p_{0}+1,p_{0}+2,\ldots,m+1+p\right\}.$ (38)
Then the remainder of the matrix exponential Taylor approximation of degree $m$,
$R^{\prime}_{m}(A)=\sum_{k=m+1}^{\infty}\frac{A^{k}}{(k+1)!},$ (39)
satisfies the bound
$\|R^{\prime}_{m}(A)\|_{1}<\frac{\alpha_{p}^{m+1}}{(m+2)!}\cdot\left(\frac{1}{1-\frac{\alpha_{p}}{m+3}}\right),$ (40)
provided that condition
$\alpha_{p}<m+3,$ (41)
is fulfilled.
Finally, regarding the approximation in (8), we set the scaling parameter to $s=0$, consistent with the approach described in [25, Sec. 3.2]. For Taylor based approximations of approximation orders $m\in\{1,2,4,8,15+\}$, we employ the efficient evaluation formulas detailed in Section 3.1. For degrees exceeding these values, we utilize the classical Paterson–Stockmeyer method.
3.3 Proposed algorithms
Taking into account the evaluation formulas from Section 3.1 and the error analysis from Section 3.2, Algorithm 2 computes the matrix exponential using the scaling and squaring algorithm and the Taylor polynomial approximation. Polynomials are efficiently evaluated by means of Paterson-Stockmeyer method or evaluation formulas (10)-(17). Error tolerance $\varepsilon$ is variable, but it will be fixed to $10^{-8}$ in the experiments.
Algorithm 2 Algorithm for Computing Matrix Exponential
1:Weight matrix $W$, error tolerance $\varepsilon$
2:Matrix exponential $e^{W}$
3:Obtain the polynomial order $m$ and the scaling parameter $s$ by using Algorithms 3 or 4
[…]
5:Evaluate the matrix polynomial $X=P_{m}(W)$ by Paterson-Stockmeyer technique or by evaluation formulas (10)-(17)
6:for $i=1$ to $s$ do
7:  $X=X\cdot X$
8:end for
9:return $X$
The algorithm is fundamentally suitable for any error tolerance $\varepsilon$ that is greater than or equal to the unit roundoff error. However, note that when employing IEEE double precision arithmetic, where $u=2^{-53}\approx 1.11\times 10^{-16}\ll 10^{-8}$, to achieve a higher level of accuracy that approaches the machine precision the algorithm parameters should be adjusted. Specifically, the accuracy is improved by selecting more terms in the Taylor approximation as shown in [22, Sec. 6].
Initially, Algorithm 2 determines the optimal parameters $m$ and $s$ by invoking Algorithm 3 or Algorithm 4. Next, the matrix $W$ is scaled following the scaling and squaring approximation to reduce its norm. Then, the matrix polynomial is evaluated either using the Paterson–Stockmeyer method or evaluation formulas (10)-(17). Finally, the exponential of the matrix $W$ is recovered carrying out the appropriate $s$ matrix products corresponding to the squaring stage.
In the implemented functions derived from Algorithms 2, 3, and 4, log2 has been used to avoid the appearance of large numbers when computing high powers of $\alpha_{p}$ and factorials in expressions like (34), especially if low precision arithmetic is used.
As said above, the algorithm for evaluating (8) is analogous to Algorithm 2, setting $s=0$ and using similar evaluation formulas to those in Section 3.1 for Taylor based approximations with m=1, 2, 4, 8, 15+m=1,\,2,\,4,\,8,\,15+ and the Paterson–Stockmeyer method for Taylor approximations of degrees higher than $15+$.
Algorithm 3 determines the polynomial order $m$ and the scaling parameter $s$ required to compute the matrix exponential of $W$ using a Taylor approximation of order $m\leq 16$ and a desired tolerance $\varepsilon$. It is assumed that the polynomials are evaluated using the Paterson–Stockmeyer method.
Algorithm 3 Algorithm for Obtaining a Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via Paterson–Stockmeyer
1:Weight matrix $W$, error tolerance $\varepsilon$
2:Polynomial order $m$ and scaling parameter $s$
3:$m=0,s=0$
4:if $\|{W}\|_{1}=0$ then
5:  return $m,s$
6:end if
7:$M=[1,2,4,6,9,12,16]$
[…]
9:$K=M./J$
10:C=[1./[2!,3!,3!,4!,5!,6!,7!,8!,10!,11!,13!,14!,17!,18!]]C=[1./[2!,3!,3!,4!,5!,6!,7!,8!,10!,11!,13!,14!,17!,18!]]
11:$finish=0$
12:$i=1$
13:while $i\leq\text{length}(M)$ and finish = 0 do
14:  $m=M_{i}$
15:  $j=J_{i}$
16:  $k=K_{i}$
17:  $p=2\cdot i-1$
18:  if $m=1$ then
19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
21:  else
22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
24:  end if
25:  if $E_{1}+E_{2}\leq\varepsilon$ then
26:   $finish=1$
27:  else
28:   $i=i+1$
29:  end if
30:end while
31:if $finish=0$ then
32:  for $i=1$ to 22 do
33:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
34:   if $s1>s$ then
35:     $s=s1$
36:   end if
37:  end for
38:end if
39:if $s>20$ then
40:  $s=20$
41:end if
42:return $m,s$
Initially, if $W$ is a null matrix, the values of $m$ and $s$ equal to 0 will be returned. Next, vector $M$ stores the possible degrees of the considered polynomials, vector $J$ is completed as the power of $W$ that must be computed for each value of $M$, and vector $K$ is filled in according to the quotient of the elements of $M$ divided by the elements of $J$.
For each possible value $m$ in $M$, vector $C$ stores two consecutive values corresponding to the coefficients of the Taylor series located at positions $m+1$ and $m+2$, that is, $1/(m+1)!$ and $1/(m+2)!$.
Next, from lines 9 to 28, an attempt is made to determine the most appropriate degree $m$ of the polynomial to use, from the point of view of accuracy and computational efficiency. Since $\frac{W^{m+1}}{(m+1)!}$ and $\frac{W^{m+2}}{(m+2)!}$ are the first two terms of the remainder for the Taylor-based approximation of order $m$ to the exponential of $W$, our objective is to obtain the value of $m$ that guarantees that:
$\frac{\|W^{m+1}\|_{1}}{(m+1)!}+\frac{\|W^{m+2}\|_{1}}{(m+2)!}\leq\varepsilon,$ (42)
where $E_{1}$ corresponds to the first summand and $E_{2}$ to the second one in the algorithm.
Notwithstanding, since $W^{m+1}$ and $W^{m+2}$ are not explicitly calculated, the 1-norm of these matrix powers is bounded using the 1-norm of the highest power of $W$ computed thus far. For example, when $m=1$ (lines 16–18), $W$ is known, but $\|W^{2}\|_{1}$ and $\|W^{3}\|_{1}$ are required; these are bounded by $\|W\|_{1}^{2}$ and $\|W\|_{1}^{3}$, respectively. Similarly, if $m=12$ (lines 19–21), $W^{4}$ has been previously calculated, and the required norms $\|W^{13}\|_{1}$ and $\|W^{14}\|_{1}$ are bounded by $\|W^{4}\|_{1}^{3}\cdot\|W\|_{1}$ and $\|W^{4}\|_{1}^{3}\cdot\|W^{2}\|_{1}$. Thus, if $m$ equals 12, $j$ and $k$ take the values 4 and 3, accordingly.
If expression (42) is not satisfied for any considered value of $m$, then $m$ is set equal to the maximum degree in vector $M$, i.e., $m=16$. Consequently, the minimum value of $s$ is obtained such that
$\max\left(\frac{1}{(m+1)!}\cdot\left\|\left(\frac{W}{2^{s}}\right)^{m+1}\right\|_{1},\frac{1}{(m+2)!}\cdot\left\|\left(\frac{W}{2^{s}}\right)^{m+2}\right\|_{1}\right)\leq\varepsilon.$ (43)
By solving for $s$ in the first and second terms of (43) independently and taking the logarithm of each, we obtain
$s\geq\max\left(\frac{1}{m+1}\cdot\log_{2}\left(\frac{\|W^{m+1}\|_{1}}{\varepsilon\cdot(m+1)!}\right),\frac{1}{m+2}\cdot\log_{2}\left(\frac{\|W^{m+2}\|_{1}}{\varepsilon\cdot(m+2)!}\right)\right),$ (44)
in accordance with lines 30 to 36 in the algorithm. Finally, to avoid overscaling, the maximum value of $s$ is set to 20 in lines 37 and 38.
Similarly, Algorithm 3 fulfills the same role as Algorithm 2, but incorporates formulas (10)-(17) for evaluating the matrix polynomials. While the structure remains consistent with the previous algorithm, the polynomial degrees differ. Furthermore, the maximum explicitly computed power of $W$ is restricted to $W^{2}$. A key distinction lies in the penultimate value of vector $C$, which corresponds to the Taylor-based approximation $T_{15+}$ defined in (17). As previously noted, this approximation is equivalent to (18), where $b_{16}$ is given in IEEE double-precision arithmetic by (20). Consequently, the penultimate element of vector $C$ is derived from the corresponding remainder term (19). For $m=15$, the indices $j$ and $k$ are set to 2 and 8, respectively. Thus, using $W^{2}$, the bounds for $\|W^{16}\|_{1}$ and $\|W^{17}\|_{1}$ (lines 20, 21, and 26) are bounded by $\|W^{2}\|_{1}^{8}$ and $\|W^{2}\|_{1}^{8}\cdot\|W\|_{1}$. The calculation of the scaling parameter $s$ follows the procedure described in Algorithm 2. However, in this case determining $s$ based on the maximum value from the first two terms of the remainder (19) ensures a more conservative selection.
Algorithm 4 Algorithm for Obtaining an Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via evaluation formulas (10)-(17)
1:Weight matrix $W$, error tolerance $\varepsilon$
2:Polynomial order $m$ and scaling parameter $s$
3:$m=0,s=0$
4:if $\|{W}\|_{1}=0$ then
5:  return $m,s$
6:end if
7:$M=[1,2,4,8,15]$
8:$J=[1,2,2,2,2]$
[…]
11:$finish=0$
12:$i=1$
13:while $i\leq\text{length}(M)$ and finish = 0 do
14:  $m=M_{i}$
15:  $j=J_{i}$
16:  $k=K_{i}$
17:  $p=2\cdot i-1$
18:  if $m=1$ then
19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
21:  else
22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}$
23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}$
24:   if $j\cdot k=m$ then
25:     $E_{1}=E_{1}\cdot\|{W}\|_{1}$
26:     $E_{2}=E_{2}\cdot\|{W}^{2}\|_{1}$
27:   else
28:     $E_{2}=E_{2}\cdot\|{W}\|_{1}$
29:   end if
30:  end if
31:  if $E_{1}+E_{2}\leq\varepsilon$ then
32:   $finish=1$
33:  else
34:   $i=i+1$
35:  end if
36:end while
37:if $finish=0$ then
38:  for $i=1$ to 22 do
39:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
40:   if $s1>s$ then
41:     $s=s1$
42:   end if
43:  end for
44:end if
45:if $s>20$ then
46:  $s=20$
47:end if
48:return $m,s$
4 Numerical results
This section compares the numerical and computational properties of the different implementations. Following the baseline method in [25], we set the error tolerance to $\varepsilon=10^{-8}$. The code is available at https://github.com/hipersc/MatrixFunctions/.
In all experiments, the normwise relative error serves as our primary metric to quantify the accuracy of each matrix exponential method relative to the “exact” solution (computed to double-precision limits, as detailed below). This error is defined as:
$\textmd{Er}(A)=\frac{\|\exp(A)-\widetilde{\exp}(A)\|_{2}}{\|\exp(A)\|_{2}},$ (45)
where $\exp(A)$ and $\widetilde{\exp}(A)$ stands for the approximated and the “exact” solution.
In this section, all computational tests were conducted on a PC equipped with an Intel Core i9-14900HX processor, 32 GB of RAM, and an NVIDIA GeForce RTX 4070 graphics card. Comparisons were performed initially in MATLAB (version 2025a) and subsequently in PyTorch.
4.1 Numerical results in MATLAB
The numerical experiments are conducted against two primary benchmarks: the Taylor-based approach described in [25] (Algorithm 1) and a custom implementation of the Paterson–Stockmeyer (P–S) scheme. This selection is predicated on the current state of deep learning frameworks such as PyTorch, where native matrix exponential routines (e.g., torch.linalg.expm) are typically optimized for fixed IEEE precision formats and do not support user-defined variable error tolerances. Since [25] serves as the established reference for matrix exponential applications within generative flow models, it constitutes the appropriate baseline for assessing improvements in this specific domain. Furthermore, the inclusion of the P–S scheme as a secondary benchmark allows for a direct quantification of the efficiency gains attributable specifically to the advanced polynomial evaluation formulas integrated into the proposed algorithm.
To facilitate the comparison, the following three implementations were developed in MATLAB:
  1. 1.
    expm_flow_ps: An implementation based on the proposed Algorithm 2, where matrix polynomials are evaluated using the Paterson–Stockmeyer technique.
  2. 2.
    expm_flow_sastre: A version identical to the previous implementation, but utilizing evaluation formulas (10)-(17) for matrix polynomials.
  3. 3.
    expm_flow: The original algorithm proposed in [25], rewritten in MATLAB to ensure a fair comparison within the same computational environment.
To evaluate the performance of these three codes, a testbed composed of 360 ill-conditioned matrices belonging to the Matrix Computation Toolbox (MCT) [35] the Eigtool MATLAB Package (EMP) [36] was generated. Their size, always a power of 2, ranged from 4 to 1024. The “exact” exponential function of each matrix $A$ was calculated, where applicable, following the three steps outlined in the following procedure:
  • 1.
    By applying the MATLAB function eig to the input matrix $A$, the matrices $D$ and $V$ were obtained such that $A=V\cdot D\cdot V^{-1}$. The matrix $E_{1}=V\cdot\exp(D)\cdot V^{-1}$ was then generated.
  • 2.
    The matrix $E_{2}=\exp(A)$ was derived using the MATLAB expm function, along with MATLAB Symbolic Math Toolbox and the function vpa (variable-precision floating-point arithmetic) with 256 decimal digits.
  • 3.
    Matrix $E_{1}$ was considered the “exact” exponential of $A$ if the following condition holds:
    $\frac{{{{\left\|{{E_{1}}-{E_{2}}}\right\|}_{2}}}}{{{{\left\|{{E_{1}}}\right\|}_{2}}}}\leq u,$
    where $u=2^{-53}$ stands for the unit roundoff in IEEE double precision arithmetic. Otherwise, matrix $A$ was excluded.
Graphically, Figure 1 shows the results obtained from the first of the experiments, carried out on the matrices comprising MCT and EMP. Figure 1(a) displays the normwise relative error committed by each of the codes for the different matrices analysed. These same errors, ordered from highest to lowest for code expm_flow_sastre, are also shown in Figure 1(b). The black line appearing in 1(a) corresponds to the expected relative error when calculating the exponential of the matrices. This value is equivalent to the product of the condition number of the exponential function for each matrix multiplied by the error tolerance used ($10^{-8}$) in our experiments. Ideally, it is to be expected that all results will be below this line, with those ones further down clearly being better. It is easy to see how the results for expm_flow are less accurate than those for expm_flow_sastre and expm_flow_ps, as they occupy the highest positions in the graph.
Figure 1(c) corresponds to the performance profile. It represents, for each value $\alpha$ on the $x$-axis, the percentage of matrices, expressed as a proportion of one, for which the error incurred in calculating their exponential function is $\alpha$ times less than or equal to the minimum error committed by each of the methods compared. The higher the values provided, the better the code. As we can appreciate, expm_flow_sastre and expm_flow_ps occupy the top of the picture, while expm_flow provides the poorest results. While expm_flow_ps offers better results for $\alpha$ equal to 1, this improvement is reversed in favour of expm_flow_sastre from $\alpha$ values close to 4.
The direct correspondence of these values for $\alpha$ equal to 1 are those ones collected, in the form of a pie chart, on the left side of Figure 1(d). It represents the percentage of cases in which each code yielded the most accurate result. Thus, we can state that expm_flow_ps provided the closest values to the exact solution in 58% of cases, expm_flow_sastre in 42% of matrices, and expm_flow in a negligible percentage close to 0. In contrast, the right side of the picture reports the percentage of matrices for which each method gave the most inaccurate result. As we can see, expm_flow accounted for 99% of the matrices, clearly being the code with the worst performance.
As whisker plots, Figures 1(e) and Figure 1(f) display respectively the order of the polynomial applied and the value of the parameter $s$ used in the scaling and squaring technique. In each box, the red centre line indicates the median, and the blue bottom and top lines indicate the 25th and 75th percentiles, correspondingly. The whiskers extend to the most extreme data points that are not considered outliers, and outliers are represented individually by the “+” marker symbol. The code expm_flow_ps usually worked with polynomials of degree 16, or 15 in the case of expm_flow_sastre, while expm_flow used polynomials of lower order, in most cases between 5 and 7. Regarding the scaling parameter $s$, the median value was 1 for the code expm_flow_ps, 2 for expm_flow_sastre, and 5 for expm_flow. The maximum values reached were equal to 20 for expm_flow_ps and expm_flow_sastre and 718 for expm_flow.
Finally, Figures 1(g) and 1(h) present, in the form of a bar chart, two key comparative results: the number of matrix products required by each code and the response times (expressed in seconds) employed in the computation of all test matrices. Specifically, expm_flow_ps required 3,110 products, which is approximately $1.20$ times the 2,597 products needed by the baseline expm_flow_sastre. The least efficient code, expm_flow, required 5,397 matrix products, representing a factor of approximately $2.08$ compared to expm_flow_sastre. This comparison highlights expm_flow_sastre as the most product-efficient implementation, followed by expm_flow_ps, with expm_flow consistently requiring over twice the computational effort in terms of matrix multiplications.
The analysis of execution times confirms the relative performance difference, using expm_flow_sastre as the baseline for comparison. Specifically, expm_flow_ps required $1.56$ seconds for computation, making it approximately $1.13$ times slower than the $1.39$ seconds required by expm_flow_sastre. The least efficient code, expm_flow, required $2.31$ seconds, resulting in a time factor of approximately $1.67$ times slower compared to expm_flow_sastre. This comparison confirms expm_flow_sastre is the fastest implementation, with expm_flow exhibiting the largest runtime overhead. Notably, these execution time ratios are lower than the corresponding matrix product ratios. This reduced proportionality occurs because the test matrices, which range in size from $4\times 4$ to $1024\times 1024$, are not all large enough for the number of computed matrix products to fully dominate the total execution cost.
(a) Normwise relative error.
(b) Ordered normwise relative error.
(c) Performance profile.
(d) Lowest and highest relative error rate.
(e) Polynomial order.
(f) Scaling parameter.
(g) Number of matrix products.
(h) Execution time.
Figure 1: Experimental results for matrices in MCT and EMP sets.
4.2 Performance evaluation in the PyTorch generative flow environment
In this subsection, the results of tests performed directly on the code available in [25] will be described. As mention before, this code incorporates a matrix exponential function into neural networks implementing generative flows models. In our executions, the matrix exponential is computed using the following functions implemented in PyTorch:
  1. 1.
    expm_flow_ps is based on Algorithm 2 and evaluates the polynomial using the Paterson–Stockmeyer method.
  2. 2.
    expm_flow_sastre: Analogous to the previous function, this implementation is based on Algorithm 2, but the polynomials are evaluated using formulas (10)-(17).
  3. 3.
    expm_flow directly corresponds to the matrix exponential approximation method presented in [25].
All executions were carried out using CUDA and the graphics card of the machine involved. As testbeds in training, CIFAR-10 [37], ImageNet32 and ImageNet64 [38] datasets were used. The number of invocations analysed to extract results to the functions responsible for calculating the matrix exponential during the training process was 5000. For each call to the exponential function, multiple data were recorded for the matrices comprising the tensor to be computed, such as the number of matrices, their sizes, and the highest matrix $\infty$-norm of all of them. Thus, we can confirm that the $\infty$-norm of the matrices studied varied from 2.84e-4 to 12.57 for CIFAR-10, from 1.17e-05 to 12.49 for ImageNet32, and from 1.27e-05 to 12.8 for ImageNet64.
In addition, the following metrics were logged for each of the three comparison functions: the degree of the polynomial used, the scaling parameter, the maximum error observed across all matrices in the tensor, the total number of matrix products involved, and the required response time.
When estimating the normwise relative error committed in the computation of each matrix in (45), it was assumed that linalg.matrix_exp, the PyTorch implementation of the exponential function, provides an exact solution.
Figures 2, 3, and 4 graphically illustrate the results obtained for CIFAR-10, ImageNet32 and ImageNet64 datasets, respectively. Figures 2(a), 3(a), and 4(a) plot the normwise relative errors just for the first 150 training runs. For clarity, these same results are shown sorted, but now for the 5000 runs, from largest to smallest error independently for the three functions analysed in Figures 2(b), 3(b), and 4(b). Unlike 1(b), there is here no correspondence between a matrix on the $x$-axis and the three errors that appear on the $y$-axis, due to the aforementioned sorting. As we can observe, expm_flow always resulted in higher errors in the exponential approximation. In contrast, expm_flow_sastre and expm_flow_ps provided very similar results, with expm_flow_sastre being slightly more accurate for the three datasets. This conclusion can also be deduced from the Figures 2(c), 3(c), and 4(c) related to the performance profile, where expm_flow_sastre and expm_flow_ps quickly provided probabilities equal to 1.
Figures2(d), 3(d), and 4(d) also confirm that code expm_flow_sastre delivered a more accurate result than the other functions in a higher percentage of cases, with values of 50%, 49%, and 50% of the matrices for the three datasets, respectively. In turn, function expm_flow_sastre provided similar but slightly lower percentages, with values of 48%, for CIFAR-10, and 45%, for ImageNet32 and ImageNet64.
In clear accordance with Figures 1(e), Figures 2(e), 3(e), and 4(e) also reveal that expm_flow used polynomials of a lower order than those of the other functions. In contrast, expm_flow required higher values of the scaling parameter, as can be appreciated in Figures 2(f), 3(f), and 4(f).
Finally, with regard to the number of products, Figures 2(g), 3(g), and 4(g) state that expm_flow carried out 1.99, 1.86, and 1.88 times more products than expm_flow_sastre in the calculation of matrix exponentials. Similarly, Figures 2(h), 3(h), and 4(h) show that the exponential computation based on function expm_flow_sastre was 1.87, 1.97, and 2.5 times faster than the same calculation using function expm_flow.
(a) Normwise relative error.
(b) Unmatched ordered normwise relative error.
(c) Performance profile.
(d) Lowest and highest relative error rate.
(e) Polynomial order.
(f) Scaling parameter.
(g) Number of matrix products.
(h) Execution time.
Figure 2: Experimental results for CIFAR10 dataset.
(a) Normwise relative error.
(b) Unmatched ordered normwise relative error.
(c) Performance profile.
(d) Lowest and highest relative error rate.
(e) Polynomial order.
(f) Scaling parameter.
(g) Number of matrix products.
(h) Execution time.
Figure 3: Experimental results for ImageNet32 dataset.
(a) Normwise relative error.
(b) Unmatched ordered normwise relative error.
(c) Performance profile.
(d) Lowest and highest relative error rate.
(e) Polynomial order.
(f) Scaling parameter.
(g) Number of matrix products.
(h) Execution time.
Figure 4: Experimental results for ImageNet64 dataset.
5 Performance analysis in generative flow models
To evaluate the efficiency gains of the proposed matrix exponential approach (expm_flow_sastre) relative to the baseline (expm_flow), we conducted a series of numerical experiments using three standard computer vision datasets: CIFAR-10, ImageNet32, and ImageNet64. All tests were executed on a system equipped with an Intel Xeon Platinum 8268 Processor, an NVIDIA RTX8000 GPU and 128 GB of RAM.
The evaluation consists of two parts. First, isolated benchmarks were performed to compare the execution times of both methods across various matrix sizes, independent of other neural network components. Second, both algorithms were integrated into the coupling layers of a Glow-based architecture [39], following the methodology described in [25]. Performance is quantified by measuring total computation time and calculating the speed-up ratio, defined as the execution time of expm_flow divided by that of expm_flow_sastre.
The flow model architecture, illustrated in Figure 5, follows a multi-scale structure. The squeeze operation performs downsampling by halving the spatial dimensions and quadrupling the channel count. Subsequently, the split operation partitions the feature maps along the channel dimension; one portion is directed to the output as a latent representation, while the remaining part is propagated through the subsequent layers of the model. Training was conducted for 50 epochs using the Adam optimizer [43] with a fixed learning rate of 0.01.
Figure 5: Diagram overview of flow model trained to generate images.
5.1 Computational scaling and matrix dimensionality
This analysis evaluates how the performance of the proposed matrix exponential implementation, expm_flow_sastre, scales relative to the baseline, expm_flow, as matrix dimensionality increases. We examine matrix orders $n\in\{8,16,32,64,128,256,512,1024\}$ under two distinct scenarios: individual matrix transformations of size $n\times n$ and batched tensor operations of size $n\times 16\times 16$. The execution time required to compute 1000 matrix exponentials was recorded for each configuration, with the resulting trends illustrated in Figure 6.
In both experimental setups, the relative efficiency of the proposed algorithm improves as the matrix order increases. A more pronounced performance gain is observed when varying the batch size compared to isolated changes in matrix dimensionality. This behavior stems from the fact that for smaller matrices, such as $8\times 8$ or $16\times 16$, the total execution cost is not yet dominated by the number of matrix products, as fixed computational overheads remain significant. As the dimensionality grows toward $1024\times 1024$, the computation becomes strictly limited by the number of matrix multiplications $M$. Consequently, the proposed method yields substantial benefits by reducing the required multiplication count, effectively leveraging the parallel linear algebra routines provided by the PyTorch backend, such as cuBLAS and MKL.
Figure 6: Comparison of the execution time required to calculate the matrix exponential using the original method (expm_flow) and the proposed method (expm_flow_sastre). The results are presented for different matrix sizes $n\in\{2,4,8,16,32,64,128,256,512,1024\}$. (Left) Performance analysis for single $n\times n$ matrices. (Right) Performance analysis for tensors of size $n\times 16\times 16$.
5.2 Training Performance and Scaling
To evaluate the practical utility of our approach, we integrated the proposed matrix exponential method into a generative framework. Glow [39] is a coupling layer approach to learning invertible transformations. We replaced the original implementation (expm_flow) with our optimized function (expm_flow_sastre) to measure its impact on computational efficiency during the training phase. All models were trained for 50 epochs across three benchmark datasets: CIFAR-10, ImageNet32, and ImageNet64. As detailed in Table 4, the proposed method demonstrates increased computational efficiency across all tested configurations. While the minimum observed speed-up was $3.91$ for ImageNet64, the most significant gain occurred with ImageNet32, where training was nearly 10 times faster ($9.74$) than the baseline. This dramatic reduction in training time, specifically the shift from 32,304 seconds to 3,317 seconds per epoch on ImageNet32, demonstrates that the computational bottleneck previously associated with matrix exponentials in flow-based models can be effectively mitigated.
Dataset CIFAR-10 ImageNet32 ImageNet64
expm_flow time 2463 32304 42482
expm_flow_sastre time 443.4 3317 10875
Speed-up 5.554 9.739 3.906
Table 4: Average training time per epoch (in seconds) for the matrix exponential version of Glow. Results compare the efficiency of our proposed expm_flow_sastre function against the original expm_flow implementation across three standard datasets.
5.3 Inference and Sampling Performance
While the previous sections evaluated performance during the training phase, efficiency during inference is equally critical for generative models. Table 5 presents a comparison of the sampling times between expm_flow and expm_flow_sastre. In this context, sampling time refers to the duration required for the model to generate an image from the latent space. We evaluated two scenarios: generating a single image ($n=1$) and generating a batch of 128 images ($n=128$). The results demonstrate that the speedup provided by our method becomes significantly more pronounced as the batch size increases. While the performance is nearly identical for a single sample, the expm_flow_sastre variant achieves a speedup factor of approximately $1.95$ for larger batches, effectively halving the inference latency.
Sample 1 sample 128 samples
expm_flow time 0.3839 0.8286
expm_flow_sastre time 0.3836 0.4247
Speed-up 1.001 1.951
Table 5: Average inference times (in seconds) for sampling a matrix exponential variant of Glow. We compare the original expm_flow against expm_flow_sastre for single and batched samples, including the calculated speedup factor.
6 Conclusions
This work presented a Taylor-based algorithm for the computation of the matrix exponential, utilizing advanced polynomial evaluation schemes that extend beyond the classical Paterson–Stockmeyer technique. The proposed method incorporates a self-contained framework for the dynamic selection of the Taylor order and scaling factor, based on a rigorous error analysis designed to maintain accuracy for any user-specified tolerance $\varepsilon$ greater than or equal to the working unit roundoff. This approach distinguishes itself from traditional implementations that rely on fixed-precision tables and from complex variable-precision algorithms that require specialized external libraries.
The algorithm was specifically evaluated within the context of generative AI flows, where the matrix exponential constitutes a primary computational requirement for modeling invertible transformations. By leveraging refined error bounds for nonnormal matrices and streamlined operational logic, the proposed method provides a more efficient alternative to previous dynamic schemes. Numerical evaluations conducted on both a diverse testbed of matrices and practical generative models demonstrate that, relative to the original implementation, the proposed Taylor-based approach yields training speedups between $3.91$ and $9.74$ while reducing inference latency by approximately 50% in large-batch scenarios.
These results demonstrate that modern polynomial evaluation strategies allow for higher-order approximations at a lower computational cost, measured in terms of matrix multiplications ($M$), than traditional rational or Paterson–Stockmeyer-based methods. Consequently, the proposed implementation offers a portable, library-independent solution for high-throughput applications and large-scale simulation environments.
Acknowledgements
The authors would like to thank Roger B. Dannenberg for providing valuable feedback on the writing of this manuscript.
Author contributions
J. Sastre: Conceptualization, Methodology, Software, Formal analysis, Funding acquisition, Project administration, Writing - original draft, Writing - review & editing.
D. Faronbi: Software, Validation, Investigation, Data curation, Resources, Visualization, Writing - original draft, Writing - review & editing.
J.M. Alonso: Methodology, Software, Formal analysis, Validation, Investigation, Data curation, Resources, Visualization, Writing - original draft, Writing - review & editing.
P. Traver: Software, Investigation, Validation, Resources, Visualization, Writing - review & editing.
J. Ibáñez: Methodology, Formal analysis, Investigation, Data curation, Visualization, Supervision, Writing - review & editing.
N. Lloret.: Resources, Validation, Supervision, Funding acquisition, Project administration, Writing - review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Disclosure of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Gemini in order to improve the language and readability of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v2

@@
 For instance, the method presented in [20] introduced $m+$ approximations, where the first $m$ terms match the Taylor series exactly, while additional terms are optimized to enhance accuracy.
-These schemes, such as the $m=15+$ case introduced in [9, Ex.
+These schemes, such as the $m=15+$ case introduced in [8, Ex.
 5.1], achieve greater precision than standard Taylor polynomials of equivalent computational cost.
-Furthermore, recent optimizations for determining evaluation formula coefficients [18] have shown promising results for broader matrix function computations [18, 19].
+Furthermore, recent optimizations for determining the coefficients of evaluation formulas [17] have shown promising results for broader matrix function computations [17, 18].
 The cost of these polynomial approximation methods is dominated by matrix multiplications, denoted by $M$.
 In practice, these operations scale as $\mathcal{O}(n^{3})$ for $n\times n$ matrices, whereas matrix additions and scalar multiplications are of order $\mathcal{O}(n^{2})$.
-In rational methods, matrix inverses are computed by solving linear systems, the cost of which can also be expressed in terms of matrix products as [12, App.
+In rational methods, matrix inverses are computed by solving linear systems, the cost of which can also be expressed in terms of matrix products as [11, App.
 C]:
@@
 Consequently, throughout this paper, the cost of both rational and polynomial approximations is quantified in terms of $M$.
-Table 1 compares the cost and achievable approximation orders of various state-of-the-art methods, including Padé approximants, Paterson–Stockmeyer evaluation, and mixed rational-polynomial schemes [11, 22, 14].
+Table 1 compares the cost and achievable approximation orders of various state-of-the-art methods, including Padé approximants [21], Paterson–Stockmeyer evaluation [12], advanced matrix polynomial evaluation [8, 20, 13] and mixed rational-polynomial schemes [10].
 The results highlight the superior efficiency of modern polynomial evaluation strategies, which facilitate higher-order approximations using fewer matrix multiplications.
-In this paper, we propose an optimized implementation of the matrix exponential based on the evaluation methods described in [9].
+In this paper, we propose an optimized implementation of the matrix exponential based on the matrix polynomial evaluation methods described in [8].
 Our approach is specifically tailored for the high-throughput requirements of Generative AI flows, where the repeated evaluation of the matrix exponential often constitutes a primary computational bottleneck.
@@
 Approx. order $m$ [20] 8 $15+$ $21+$ 24 30
+Padé evaluation cost $3.33M$ $4.33M$ $5.33M$ $6.33M$ $7.33M$
+Approx. order Padé method [21, Tab. 2.2] 6 10 14 18 26
 Mixed rational polynomial approx. cost $3.33M$ $4.33M$ $5.33M$ $6M$ $7M$
-Approx. order from method [11, Tab. 3] 9 12 16 21 28
-Padé evaluation cost $3.33M$ $4.33M$ $5.33M$ $6.33M$ $7.33M$
-Approx. order Padé method [23, Tab. 2.2] 6 10 14 18 26
-Table 1: Cost and available order of approximation with the polynomial evaluation methods from Paterson–Stockmeyer [13], Sastre–Ibáñez–Defez [22, Sec.
-3], Bader–Blanes–Casas [14, Sec.
-4], the mixed rational and polynomial approximation method from Sastre [11, Tab.
-3] and the Padé approximation method from Higham [23, Tab.
-2.2].
+Approx. order from method [10, Tab. 3] 9 12 16 21 28
+Table 1: Cost and available order of approximation with the polynomial evaluation methods from Paterson–Stockmeyer [12], Bader–Blanes–Casas [13, Sec.
+4], Sastre–Ibáñez–Defez [20, Sec.
+3], the Padé approximation method from Higham [21, Tab.
+2.2] and the mixed rational and polynomial approximation method from Sastre [10, Tab.
+3].
 Throughout this manuscript, $\mathbb{C}^{n\times n}$ denotes the set of complex matrices of size $n\times n$, $I$ denotes the identity matrix for this set, $\rho(A)$ is the spectral radius of matrix $A$, and $\mathbb{N}$ denotes the set of positive integers.
@@
 This guarantees invertibility, improves expressiveness over triangular matrices, and enables efficient computation of both the inverse and the log-determinant.
-Specifically, the inverse is given by $(e^{W})^{-1}=e^{-W}$, and the log-determinant simplifies to $\log\det(e^{W})=\operatorname{Tr}(W)$, reducing the computational cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
+Specifically, the inverse is given by $(e^{W})^{-1}=e^{-W}$, and the log-determinant simplifies to $\log|(e^{W})|=\operatorname{Tr}(W)$, reducing the computational cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
 The matrix exponential maps $\mathbb{M}_{n}(\mathbb{R})$ to $\mathrm{GL}_{n}(\mathbb{R})^{+}$, the group of invertible matrices with positive determinant.
@@
 16:return $Y$
-For the sake of completeness the algorithm for the computation of the matrix exponential described in [25, Sec.
+For the sake of completeness the algorithm for the computation of the matrix exponential described in [23, Sec.
 3.2] is reproduced here as Algorithm 1.
@@
 $\|R_{m}(W)\|_{1}\leq\frac{\|W\|_{1}^{m+1}}{(m+1)!}\left(\frac{1}{1-\frac{\|W\|_{1}}{m+2}}\right).$ (6)
-Note that the truncated Taylor series is particularly well-suited for scenarios with limited computational resources, restricted access to external libraries or large-scale problems involving well-behaved matrices [26, p.
+Note that the truncated Taylor series is particularly well-suited for scenarios with limited computational resources, restricted access to external libraries or large-scale problems involving well-behaved matrices [24, p.
 52].
-According to [25, Sec.
-3.2], the weight matrices in neural networks, denoted by $W$, are often small in norm such that $|W|_{1}$ is also small.
+According to [23, Sec.
+3.2], the weight matrices in neural networks, denoted by $W$, are often small in norm such that $\|W\|_{1}$ is also small.
 This property makes the incorporation of the matrix exponential into neural network architectures computationally feasible.
 Exploiting the identity $e^{W}=(e^{W/2^{s}})^{2^{s}}$, in Algorithm 1 the matrix is first scaled to ensure that $|W/2^{s}|<1/2$, after which the exponential is computed and subsequently squared $2^{s}$ times.
-This scaling and squaring strategy further reduces the computational cost while maintaining numerical stability; see [25, Alg.
+This scaling and squaring strategy further reduces the computational cost while maintaining numerical stability; see [23, Alg.
 1].
-The computational cost of evaluating a Taylor approximation of degree $m$ using [25, Alg.
-1], denoted by $T_{m}(A)$, is measured in terms of matrix multiplications.
-Let $C_{\text{orig}}$ represent the cost of evaluating the unscaled Taylor polynomial in Algorithm [25, Alg.
+The computational cost of evaluating the Taylor approximation of degree $m$ from (4) using [23, Alg.
+1] is measured in terms of matrix multiplications $(M)$.
+Let $C_{\text{orig}}$ represent the cost of evaluating the unscaled Taylor polynomial in Algorithm [23, Alg.
 1], which is given by
-$C_{\text{orig}}=m-1.$ (7)
+$C_{\text{orig}}=(m-1)M.$ (7)
 When scaling is applied, the total cost becomes (s+m1)M(s+m-1)M, where $s$ is the scaling parameter.
-As reported in [25, Sec.
+As reported in [23, Sec.
 6.4], a tolerance of $\varepsilon=10^{-8}$ is used.
-Under this setting, the total number of matrix products, given by s+m1s+m-1, does not exceed 11, with an average value of 9.28 observed during training [25, Tab.
+Under this setting, the total number of matrix products, given by s+m1s+m-1, does not exceed 11, with an average value of 9.28 observed during training [23, Tab.
 6].
-Cost (7) makes Algorithm 1 unsuitable for high-dimensional data [25, Sec.
+Cost (7) makes Algorithm 1 unsuitable for high-dimensional data [23, Sec.
 3.2].
@@
 $e^{W}\approx I+A_{1}\left(\sum_{i=0}^{m}\frac{V^{i}}{(i+1)!}\right)A_{2},$ (8)
-where the truncation error bound of the matrix series of $V$, denoted here by $R^{\prime}_{m}(V)$, is given by (8) from [25, Sec. 3.2] reproduced here
+where the truncation error bound of the matrix series of $V$, denoted here by $R^{\prime}_{m}(V)$, is given by (8) from [23, Sec. 3.2] reproduced here
 $\left\|R^{\prime}_{m}(V)\right\|_{1}=\left\|\sum_{i=m+1}^{\infty}\frac{V^{i}}{(i+1)!}\right\|_{1}\leq\left(\frac{\|V\|_{1}^{m+1}}{(m+2)!}\right)\left(\frac{1}{1-\frac{\|V\|_{1}}{m+3}}\right).$ (9)
@@
 Thus, there is a trade-off between computational cost and model capacity.
-The approximation method is similar to Algorithm 1, with modifications: set the scale coefficient $s:=0$, change line 4 to $Y:=W/2$, and line 5 to $k:=3$ [25, Sec. 3.2].
-While the cost of the algorithms proposed in [25, Sec. 3.2] represent a reasonable computational cost, further reductions are possible by optimizing the polynomial evaluation strategy.
+The approximation method is similar to Algorithm 1, with modifications: set the scale coefficient $s:=0$, change line 4 to $Y:=W/2$, and line 5 to $k:=3$ [23, Sec. 3.2].
+While the cost of the algorithms proposed in [23, Sec. 3.2] represent a reasonable computational cost, further reductions are possible by optimizing the polynomial evaluation strategy.
 3 Proposed method
@@
 Furthermore, the algorithm incorporates a dynamic procedure that optimally selects the Taylor series degree ($m$) and the scaling factor ($s$) to minimize the overall computational cost.
-This selection ensures that the error tolerance is strictly maintained, chosen to be no less than the unit roundoff ($u$) of the arithmetic used, while also strategically reducing the risk of overscaling [21, Sec.
+This selection ensures that the error tolerance is maintained, chosen to be no less than the unit roundoff ($u$) of the arithmetic used, while also strategically reducing the risk of overscaling [19, Sec.
 1].
@@
     The cost of evaluating $T_{8}(A)$ using formulas (13) and (3.1) is $3M$: one multiplication for computing $A^{2}$, one for evaluating (13), and one for evaluating (3.1).
-In contrast, computing the Taylor approximation of degree $m=8$ using the original algorithm from [25, Alg.
+In contrast, computing the Taylor approximation of degree $m=8$ using the original algorithm from [23, Alg.
 1], as given by (7), requires $7M$, which is more than twice as expensive.
@@
 Approximations of this type are referred to in [20] as $T_{15+}(A)$.
-Similarly, in this case, the cost of evaluating $y_{22}(A)$ using formulas (15)-(17) is $4M$.
+Similarly, the evaluation of $y_{22}(A)$ via formulas (15)–(17) requires only $4M$ matrix products.
+This cost contrast is even more pronounced when using the low-rank parameterization from (8).
+Since no scaling is applied in that case, the evaluation of the corresponding matrix polynomial of degree 15 requires $14M$ products.
+Consequently, the baseline approach is 3.5 times more expensive than our proposed method in this configuration.
 3.2 Error analysis
@@
 $\|A^{k}\|\leq\|A\|^{k},\quad k=1,2,\dots,\infty,$ (22)
-can be significantly strict, i.e., $\|A^{k}\|\ll\|A\|^{k}$, as discussed in [21, Sec.
+can be significantly strict, i.e., $\|A^{k}\|\ll\|A\|^{k}$, as discussed in [19, Sec.
 1].
@@
 Proof.
-Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. The remainder of the matrix exponential Taylor approximation of degree $m$ given by (5)
+Let $A\in\mathbb{C}^{n\times n}$ be a square matrix and let $m\in\mathbb{N}$. The remainder of the matrix exponential Taylor approximation of degree $m$ is given by (5)
 $R_{m}(A)=e^{A}-T_{m}(A)=\sum_{k=m+1}^{\infty}\frac{A^{k}}{k!}.$
@@
 $(m+2)(m+3)\cdots(m+1+k)>(m+2)^{k},$
-so
+so for $x>0$ it follows that
 $\frac{x^{k}}{(m+2)(m+3)\cdots(m+1+k)}<\left(\frac{x}{m+2}\right)^{k}.$
@@
 Hence,
-$\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\frac{1}{1-\frac{x}{m+2}},\quad\text{for }x<m+2.$ (30)
+$\sum_{k=m+1}^{\infty}\frac{x^{k}}{k!}=\tilde{h}_{m+1}(x)<\frac{x^{m+1}}{(m+1)!}\cdot\frac{1}{1-\frac{x}{m+2}},\quad\text{for }0<x<m+2.$ (30)
 Applying (29) and (30) with $x=\alpha_{p}<m+2$ then (27) holds. ∎
@@
 $\|R_{m}(W/2^{s})\|_{1}\leq\varepsilon\frac{1}{1-\frac{\varepsilon^{1/{m+1}}}{m+2}}.$ (37)
-We verify the fulfillment of condition (28) by utilizing the error tolerance $\varepsilon=10^{-8}$ reported in [25, Sec.
+We verify the fulfillment of condition (28) by utilizing the error tolerance $\varepsilon=10^{-8}$ reported in [23, Sec.
 6.4] and the selected Taylor degrees $m\in\{1,2,4,8,15\}$ (as defined in Section 3.1).
@@
 For the case $m=15+$, where the term $m+1=16$ is not exactly $1/16!$, selecting the maximum scaling parameter obtained with $m+1$ and also the term $m+2=17$, for which the exponential remainder term corresponds effectively to $1/17!$, will ensure a proper scaling parameter value.
-The error analysis for (8) is analogous to the error analysis described here, and the following theorem can be proved similarly:
+The error analysis for the series in (8) is analogous to the error analysis described here, and the following theorem can be proved similarly:
 Theorem 3.
@@
 is fulfilled.
-Finally, regarding the approximation in (8), we set the scaling parameter to $s=0$, consistent with the approach described in [25, Sec.
+Remark 1.
+The proposed approach offers computational advantages for nonnormal matrices. Consider the well-known test matrix (1.2) from [19], defined as
+$W=\begin{bmatrix}1&b\\ 0&-1\end{bmatrix}$,
+taking for instance $b=10$.
+For this matrix, the baseline Algorithm 1 selects a scaling parameter $s=5$, whereas it is easy to show that our approach, based on Theorem 2, yields $s=0$.
+This reduction eliminates 5 matrix-matrix multiplications, demonstrating that our method effectively prevents the over-scaling typically induced by large off-diagonal entries.
+Furthermore, unlike the approach in [19], this is achieved without estimating norms of matrix powers.
+Such estimations incur an $\mathcal{O}(n^{2})$ overhead that remains non-negligible for moderate-sized matrices, as noted in [13] for $n\leq 150$ and [20, p.
+213] for $n=128$.
+Finally, regarding the approximation in (8), we set the scaling parameter to $s=0$, consistent with the approach described in [23, Sec.
 3.2].
@@
 Taking into account the evaluation formulas from Section 3.1 and the error analysis from Section 3.2, Algorithm 2 computes the matrix exponential using the scaling and squaring algorithm and the Taylor polynomial approximation.
-Polynomials are efficiently evaluated by means of Paterson-Stockmeyer method or evaluation formulas (10)-(17).
-Error tolerance $\varepsilon$ is variable, but it will be fixed to $10^{-8}$ in the experiments.
+Polynomials are efficiently evaluated by means of Paterson–Stockmeyer method or evaluation formulas (10)-(17).
+Error tolerance $\varepsilon$ is variable, but it will be fixed to $10^{-8}$ in the experiments [23, Sec.
+6.4].
 Algorithm 2 Algorithm for Computing Matrix Exponential
@@
 […]
-5:Evaluate the matrix polynomial $X=P_{m}(W)$ by Paterson-Stockmeyer technique or by evaluation formulas (10)-(17)
+5:Evaluate the matrix polynomial $X=P_{m}(W)$ by Paterson–Stockmeyer technique or by evaluation formulas (10)-(17)
 6:for $i=1$ to $s$ do
@@
 However, note that when employing IEEE double precision arithmetic, where $u=2^{-53}\approx 1.11\times 10^{-16}\ll 10^{-8}$, to achieve a higher level of accuracy that approaches the machine precision the algorithm parameters should be adjusted.
-Specifically, the accuracy is improved by selecting more terms in the Taylor approximation as shown in [22, Sec.
+Specifically, the accuracy is improved by selecting more terms in the Taylor approximation as shown in [20, Sec.
 6].
-Initially, Algorithm 2 determines the optimal parameters $m$ and $s$ by invoking Algorithm 3 or Algorithm 4.
+Initially, Algorithm 2 determines the optimal parameters $m$ and $s$ by invoking Algorithm 3.3 or Algorithm 3.3.
 Next, the matrix $W$ is scaled following the scaling and squaring approximation to reduce its norm.
@@
 Finally, the exponential of the matrix $W$ is recovered carrying out the appropriate $s$ matrix products corresponding to the squaring stage.
-In the implemented functions derived from Algorithms 2, 3, and 4, log2 has been used to avoid the appearance of large numbers when computing high powers of $\alpha_{p}$ and factorials in expressions like (34), especially if low precision arithmetic is used.
+In the implemented functions derived from Algorithms 2, 3.3, and 3.3, log2 has been used to avoid the appearance of large numbers when computing high powers of $\alpha_{p}$ and factorials in expressions like (34), especially if low precision arithmetic is used.
 As said above, the algorithm for evaluating (8) is analogous to Algorithm 2, setting $s=0$ and using similar evaluation formulas to those in Section 3.1 for Taylor based approximations with m=1, 2, 4, 8, 15+m=1,\,2,\,4,\,8,\,15+ and the Paterson–Stockmeyer method for Taylor approximations of degrees higher than $15+$.
-Algorithm 3 determines the polynomial order $m$ and the scaling parameter $s$ required to compute the matrix exponential of $W$ using a Taylor approximation of order $m\leq 16$ and a desired tolerance $\varepsilon$.
+Algorithm 3.3 determines the polynomial order $m$ and the scaling parameter $s$ required to compute the matrix exponential of $W$ using a Taylor approximation of order $m\leq 16$ and a desired tolerance $\varepsilon$.
 It is assumed that the polynomials are evaluated using the Paterson–Stockmeyer method.
@@
 16:  $k=K_{i}$
-17:  $p=2\cdot i-1$
-18:  if $m=1$ then
-19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
-20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
-21:  else
-22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
-23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
-24:  end if
-25:  if $E_{1}+E_{2}\leq\varepsilon$ then
-26:   $finish=1$
-27:  else
-28:   $i=i+1$
-29:  end if
-30:end while
-31:if $finish=0$ then
-32:  for $i=1$ to 22 do
-33:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
-34:   if $s1>s$ then
-35:     $s=s1$
-36:   end if
-37:  end for
-38:end if
-39:if $s>20$ then
-40:  $s=20$
+17:  if $j>k$ then
+18:    $W^{j}=W^{j-1}\cdot W$
+19:  end if
+20:  $p=2\cdot i-1$
+21:  if $m=1$ then
+22:    $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
+23:    $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
+24:  else
+25:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
+26:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
+27:  end if
+28:  if $E_{1}+E_{2}\leq\varepsilon$ then
+29:    $finish=1$
+30:  else
+31:    $i=i+1$
+32:  end if
+33:end while
+34:if $finish=0$ then
+35:  for $i=1$ to 22 do
+36:    $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
+37:    if $s1>s$ then
+38:     $s=s1$
+39:    end if
+40:  end for
 41:end if
-42:return $m,s$
+42:if $s>20$ then
+43:  $s=20$
+44:end if
+45:return $m,s$
 Initially, if $W$ is a null matrix, the values of $m$ and $s$ equal to 0 will be returned.
@@
 For each possible value $m$ in $M$, vector $C$ stores two consecutive values corresponding to the coefficients of the Taylor series located at positions $m+1$ and $m+2$, that is, $1/(m+1)!$ and $1/(m+2)!$.
-Next, from lines 9 to 28, an attempt is made to determine the most appropriate degree $m$ of the polynomial to use, from the point of view of accuracy and computational efficiency.
+Next, from lines 9 to 31, an attempt is made to determine the most appropriate degree $m$ of the polynomial to use, from the point of view of accuracy and computational efficiency.
 Since $\frac{W^{m+1}}{(m+1)!}$ and $\frac{W^{m+2}}{(m+2)!}$ are the first two terms of the remainder for the Taylor-based approximation of order $m$ to the exponential of $W$, our objective is to obtain the value of $m$ that guarantees that:
@@
 Notwithstanding, since $W^{m+1}$ and $W^{m+2}$ are not explicitly calculated, the 1-norm of these matrix powers is bounded using the 1-norm of the highest power of $W$ computed thus far.
-For example, when $m=1$ (lines 16–18), $W$ is known, but $\|W^{2}\|_{1}$ and $\|W^{3}\|_{1}$ are required; these are bounded by $\|W\|_{1}^{2}$ and $\|W\|_{1}^{3}$, respectively.
-Similarly, if $m=12$ (lines 19–21), $W^{4}$ has been previously calculated, and the required norms $\|W^{13}\|_{1}$ and $\|W^{14}\|_{1}$ are bounded by $\|W^{4}\|_{1}^{3}\cdot\|W\|_{1}$ and $\|W^{4}\|_{1}^{3}\cdot\|W^{2}\|_{1}$.
+For example, when $m=1$ (lines 19–21), just $W$ is known, but $\|W^{2}\|_{1}$ and $\|W^{3}\|_{1}$ are required; these are bounded by $\|W\|_{1}^{2}$ and $\|W\|_{1}^{3}$, respectively.
+Similarly, if, for example, $m=12$ (lines 22–24), $W^{4}$ will have been previously calculated, and the required norms $\|W^{13}\|_{1}$ and $\|W^{14}\|_{1}$ are bounded by $\|W^{4}\|_{1}^{3}\cdot\|W\|_{1}$ and $\|W^{4}\|_{1}^{3}\cdot\|W^{2}\|_{1}$.
 Thus, if $m$ equals 12, $j$ and $k$ take the values 4 and 3, accordingly.
+In detail, Table 4 contains the values of $j$ and $k$ used for any value of $m$ taken into account.
 If expression (42) is not satisfied for any considered value of $m$, then $m$ is set equal to the maximum degree in vector $M$, i.e., $m=16$.
@@
 $s\geq\max\left(\frac{1}{m+1}\cdot\log_{2}\left(\frac{\|W^{m+1}\|_{1}}{\varepsilon\cdot(m+1)!}\right),\frac{1}{m+2}\cdot\log_{2}\left(\frac{\|W^{m+2}\|_{1}}{\varepsilon\cdot(m+2)!}\right)\right),$ (44)
-in accordance with lines 30 to 36 in the algorithm. Finally, to avoid overscaling, the maximum value of $s$ is set to 20 in lines 37 and 38.
-Similarly, Algorithm 3 fulfills the same role as Algorithm 2, but incorporates formulas (10)-(17) for evaluating the matrix polynomials.
+in accordance with lines 32 to 39 in the algorithm. Finally, to avoid overscaling, the maximum value of $s$ is set to 20 in lines 40 and 41.
+$M$ Computed $W$ Power Required Norms Estimated Norms
+1 $W$ $\|W^{2}\|_{1}$ $\|W^{3}\|_{1}$ $\|W\|^{2}_{1}$ $\|W\|^{3}_{1}$
+2 $W^{2}$ $\|W^{3}\|_{1}$ $\|W^{4}\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W^{2}\|_{1}$
+4 $W^{2}$ $\|W^{5}\|_{1}$ $\|W^{6}\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
+6 $W^{3}$ $\|W^{7}\|_{1}$ $\|W^{8}\|_{1}$ $\|W^{3}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{3}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
+9 $W^{3}$ $\|W^{10}\|_{1}$ $\|W^{11}\|_{1}$ $\|W^{3}\|^{3}_{1}\cdot\|W\|_{1}$ $\|W^{3}\|^{3}_{1}\cdot\|W^{2}\|_{1}$
+12 $W^{4}$ $\|W^{13}\|_{1}$ $\|W^{14}\|_{1}$ $\|W^{4}\|^{3}_{1}\cdot\|W\|_{1}$ $\|W^{4}\|^{3}_{1}\cdot\|W^{2}\|_{1}$
+16 $W^{4}$ $\|W^{17}\|_{1}$ $\|W^{18}\|_{1}$ $\|W^{4}\|^{4}_{1}\cdot\|W\|_{1}$ $\|W^{4}\|^{4}_{1}\cdot\|W^{2}\|_{1}$
+$m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W^{2}\|_{1}$
+Table 4: Computed powers of $W$, the corresponding norms required by (42), and their resulting approximations for each polynomial degree $M$ considered in Algorithm 3.3.
+Similarly, Algorithm 3.3 fulfills the same role as Algorithm 3.3, but incorporates formulas (10)-(17) for evaluating the matrix polynomials.
 While the structure remains consistent with the previous algorithm, the polynomial degrees differ.
@@
 Consequently, the penultimate element of vector $C$ is derived from the corresponding remainder term (19).
+Table 5 incorporates the values of $j$ and $k$ applied for any value of $m$.
 For $m=15$, the indices $j$ and $k$ are set to 2 and 8, respectively.
-Thus, using $W^{2}$, the bounds for $\|W^{16}\|_{1}$ and $\|W^{17}\|_{1}$ (lines 20, 21, and 26) are bounded by $\|W^{2}\|_{1}^{8}$ and $\|W^{2}\|_{1}^{8}\cdot\|W\|_{1}$.
-The calculation of the scaling parameter $s$ follows the procedure described in Algorithm 2.
-However, in this case determining $s$ based on the maximum value from the first two terms of the remainder (19) ensures a more conservative selection.
-Algorithm 4 Algorithm for Obtaining an Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via evaluation formulas (10)-(17)
+Thus, using $W^{2}$, the bounds for $\|W^{16}\|_{1}$ and $\|W^{17}\|_{1}$ (lines 22, 23, and 28) are bounded by $\|W^{2}\|_{1}^{8}$ and $\|W^{2}\|_{1}^{8}\cdot\|W\|_{1}$.
+The calculation of the scaling parameter $s$ follows the procedure described in Algorithm 3.3.
+However, determining $s$ based on the maximum value from the first two terms of the remainder (19) ensures in this case a more conservative selection.
+Algorithm 4 Algorithm for Obtaining a Polynomial Order $m$ and a Scaling Parameter $s$. The polynomial will be evaluated via formulas (10)–(17)
 1:Weight matrix $W$, error tolerance $\varepsilon$
@@
 16:  $k=K_{i}$
-17:  $p=2\cdot i-1$
-18:  if $m=1$ then
-19:   $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
-20:   $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
-21:  else
-22:   $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}$
-23:   $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}$
-24:   if $j\cdot k=m$ then
-25:     $E_{1}=E_{1}\cdot\|{W}\|_{1}$
-26:     $E_{2}=E_{2}\cdot\|{W}^{2}\|_{1}$
-27:   else
-28:     $E_{2}=E_{2}\cdot\|{W}\|_{1}$
+17:  if $m=2$ then $W^{2}=W^{j-1}\cdot W$
+18:  end if
+19:  $p=2\cdot i-1$
+20:  if $m=1$ then
+21:    $E_{1}=C_{p}\cdot\|{W}\|_{1}^{2}$
+22:    $E_{2}=C_{p+1}\cdot\|{W}\|_{1}^{3}$
+23:  else if $j\cdot k=m$ then % m=2, 4, 8
+24:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
+25:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}^{2}\|_{1}$
+26:  else % m=15
+27:    $E_{1}=C_{p}\cdot\|{W^{j}}\|_{1}^{k}$
+28:    $E_{2}=C_{p+1}\cdot\|{W^{j}}\|_{1}^{k}\cdot\|{W}\|_{1}$
 29:  end if
-30:  end if
-31:  if $E_{1}+E_{2}\leq\varepsilon$ then
-32:   $finish=1$
-33:  else
-34:   $i=i+1$
-35:  end if
-36:end while
-37:if $finish=0$ then
-38:  for $i=1$ to 22 do
-39:   $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
-40:   if $s1>s$ then
-41:     $s=s1$
-42:   end if
-43:  end for
-44:end if
-45:if $s>20$ then
-46:  $s=20$
-47:end if
-48:return $m,s$
+30:  if $E_{1}+E_{2}\leq\varepsilon$ then
+31:    $finish=1$
+32:  else
+33:    $i=i+1$
+34:  end if
+35:end while
+36:if $finish=0$ then
+37:  for $i=1$ to 22 do
+38:    $s1=\left\lceil{\log_{2}(E_{i}/\varepsilon)/(m+i)}\right\rceil$
+39:    if $s1>s$ then
+40:     $s=s1$
+41:    end if
+42:  end for
+43:end if
+44:if $s>20$ then
+45:  $s=20$
+46:end if
+47:return $m,s$
+$M$ Computed $W$ Power Required Norms Estimated Norms
+1 $W$ $\|W^{2}\|_{1}$ $\|W^{3}\|_{1}$ $\|W\|^{2}_{1}$ $\|W\|^{3}_{1}$
+2 $W^{2}$ $\|W^{3}\|_{1}$ $\|W^{4}\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{1}_{1}\cdot\|W^{2}\|_{1}$
+4 $W^{2}$ $\|W^{5}\|_{1}$ $\|W^{6}\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{2}_{1}\cdot\|W^{2}\|_{1}$
+8 $W^{2}$ $\|W^{9}\|_{1}$ $\|W^{10}\|_{1}$ $\|W^{2}\|^{4}_{1}\cdot\|W\|_{1}$ $\|W^{2}\|^{4}_{1}\cdot\|W^{2}\|_{1}$
+even $m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W^{2}\|_{1}$
+15 $W^{2}$ $\|W^{16}\|_{1}$ $\|W^{17}\|_{1}$ $\|W^{2}\|^{8}_{1}$ $\|W^{2}\|^{8}_{1}\cdot\|W\|_{1}$
+odd $m$ $W^{j}$ $\|W^{m+1}\|_{1}$ $\|W^{m+2}\|_{1}$ $\|W^{j}\|^{k}_{1}$ $\|W^{j}\|^{k}_{1}\cdot\|W\|_{1}$
+Table 5: Computed powers of $W$, the norms required by (42), and their corresponding approximations for each polynomial degree $M$ considered in Algorithm 3.3. Note that $j=2$ for all degrees $m>1$.
 4 Numerical results
@@
 4.1 Numerical results in MATLAB
-The numerical experiments are conducted against two primary benchmarks: the Taylor-based approach described in [25] (Algorithm 1) and a custom implementation of the Paterson–Stockmeyer (P–S) scheme.
-This selection is predicated on the current state of deep learning frameworks such as PyTorch, where native matrix exponential routines (e.g., torch.linalg.expm) are typically optimized for fixed IEEE precision formats and do not support user-defined variable error tolerances.
-Since [25] serves as the established reference for matrix exponential applications within generative flow models, it constitutes the appropriate baseline for assessing improvements in this specific domain.
-Furthermore, the inclusion of the P–S scheme as a secondary benchmark allows for a direct quantification of the efficiency gains attributable specifically to the advanced polynomial evaluation formulas integrated into the proposed algorithm.
+The numerical experiments are evaluated against two primary benchmarks: the Taylor-based method described in [23] (Algorithm 1) and a custom implementation of the Paterson–Stockmeyer (P–S) scheme.
+This selection is predicated on the current state of deep learning frameworks such as PyTorch, where native matrix exponential routines (e.g., torch.linalg.expm) are typically optimized for standard IEEE precision formats and do not support user-defined error tolerances.
+To evaluate the performance of our proposed methods in a realistic setting, we adopt the framework from [23] as a baseline for matrix exponential applications in generative flow models.
+Furthermore, the inclusion of the P–S scheme as a secondary benchmark allows for a direct quantification of the efficiency gains attributable specifically to the optimized polynomial evaluation formulas integrated into the proposed algorithm.
 To facilitate the comparison, the following three implementations were developed in MATLAB:
   1. 1.
-    expm_flow_ps: An implementation based on the proposed Algorithm 2, where matrix polynomials are evaluated using the Paterson–Stockmeyer technique.
+    expm_flow_ps: An implementation based on the proposed Algorithms  2, where matrix polynomials are evaluated using the Paterson–Stockmeyer technique, and $3.$
   2. 2.
-    expm_flow_sastre: A version identical to the previous implementation, but utilizing evaluation formulas (10)-(17) for matrix polynomials.
+    expm_flow_opt15: A version identical to the previous implementation, but utilizing evaluation formulas (10)-(17) for matrix polynomials. Algorithm 3.3 is now considered.
   3. 3.
@@
 Graphically, Figure 1 shows the results obtained from the first of the experiments, carried out on the matrices comprising MCT and EMP.
-Figure 1(a) displays the normwise relative error committed by each of the codes for the different matrices analysed.
-These same errors, ordered from highest to lowest for code expm_flow_sastre, are also shown in Figure 1(b).
+Figure 1(a) displays the normwise relative error committed by each of the codes for the different matrices analyzed.
+These same errors, ordered from highest to lowest for code expm_flow_opt15, are also shown in Figure 1(b).
 The black line appearing in 1(a) corresponds to the expected relative error when calculating the exponential of the matrices.
@@
 Ideally, it is to be expected that all results will be below this line, with those ones further down clearly being better.
-It is easy to see how the results for expm_flow are less accurate than those for expm_flow_sastre and expm_flow_ps, as they occupy the highest positions in the graph.
+It is easy to see how the results for expm_flow are less accurate than those for expm_flow_opt15 and expm_flow_ps, as they occupy the highest positions in the graph.
 Figure 1(c) corresponds to the performance profile.
@@
 The higher the values provided, the better the code.
-As we can appreciate, expm_flow_sastre and expm_flow_ps occupy the top of the picture, while expm_flow provides the poorest results.
-While expm_flow_ps offers better results for $\alpha$ equal to 1, this improvement is reversed in favour of expm_flow_sastre from $\alpha$ values close to 4.
+As we can appreciate, expm_flow_opt15 and expm_flow_ps occupy the top of the picture, while expm_flow provides the poorest results.
+While expm_flow_ps offers better results for $\alpha$ equal to 1, this improvement is reversed in favour of expm_flow_opt15 from $\alpha$ values close to 4.
 The direct correspondence of these values for $\alpha$ equal to 1 are those ones collected, in the form of a pie chart, on the left side of Figure 1(d).
 It represents the percentage of cases in which each code yielded the most accurate result.
-Thus, we can state that expm_flow_ps provided the closest values to the exact solution in 58% of cases, expm_flow_sastre in 42% of matrices, and expm_flow in a negligible percentage close to 0.
+Thus, we can state that expm_flow_ps provided the closest values to the exact solution in 58% of cases, expm_flow_opt15 in 42% of matrices, and expm_flow in a negligible percentage close to 0.
 In contrast, the right side of the picture reports the percentage of matrices for which each method gave the most inaccurate result.
@@
 The whiskers extend to the most extreme data points that are not considered outliers, and outliers are represented individually by the “+” marker symbol.
-The code expm_flow_ps usually worked with polynomials of degree 16, or 15 in the case of expm_flow_sastre, while expm_flow used polynomials of lower order, in most cases between 5 and 7.
-Regarding the scaling parameter $s$, the median value was 1 for the code expm_flow_ps, 2 for expm_flow_sastre, and 5 for expm_flow.
-The maximum values reached were equal to 20 for expm_flow_ps and expm_flow_sastre and 718 for expm_flow.
+The code expm_flow_ps usually worked with polynomials of degree 16, or 15 in the case of expm_flow_opt15, while expm_flow used polynomials of lower order, in most cases between 5 and 7.
+Regarding the scaling parameter $s$, the median value was 1 for the code expm_flow_ps, 2 for expm_flow_opt15, and 5 for expm_flow.
+The maximum values reached were equal to 20 for expm_flow_ps and expm_flow_opt15 and 718 for expm_flow.
 Finally, Figures 1(g) and 1(h) present, in the form of a bar chart, two key comparative results: the number of matrix products required by each code and the response times (expressed in seconds) employed in the computation of all test matrices.
-Specifically, expm_flow_ps required 3,110 products, which is approximately $1.20$ times the 2,597 products needed by the baseline expm_flow_sastre.
-The least efficient code, expm_flow, required 5,397 matrix products, representing a factor of approximately $2.08$ compared to expm_flow_sastre.
-This comparison highlights expm_flow_sastre as the most product-efficient implementation, followed by expm_flow_ps, with expm_flow consistently requiring over twice the computational effort in terms of matrix multiplications.
-The analysis of execution times confirms the relative performance difference, using expm_flow_sastre as the baseline for comparison.
-Specifically, expm_flow_ps required $1.56$ seconds for computation, making it approximately $1.13$ times slower than the $1.39$ seconds required by expm_flow_sastre.
-The least efficient code, expm_flow, required $2.31$ seconds, resulting in a time factor of approximately $1.67$ times slower compared to expm_flow_sastre.
-This comparison confirms expm_flow_sastre is the fastest implementation, with expm_flow exhibiting the largest runtime overhead.
+Specifically, expm_flow_ps required 3110 products, which is approximately $1.20$ times the 2597 products needed by the baseline expm_flow_opt15.
+The least efficient code, expm_flow, required 5397 matrix products, representing a factor of approximately $2.08$ compared to expm_flow_opt15.
+This comparison highlights expm_flow_opt15 as the most product-efficient implementation, followed by expm_flow_ps, with expm_flow consistently requiring over twice the computational effort in terms of matrix multiplications.
+The analysis of execution times confirms the relative performance difference, using expm_flow_opt15 as the baseline for comparison.
+Specifically, expm_flow_ps required $1.56$ seconds for computation, making it approximately $1.13$ times slower than the $1.39$ seconds required by expm_flow_opt15.
+The least efficient code, expm_flow, required $2.31$ seconds, resulting in a time factor of approximately $1.67$ times slower compared to expm_flow_opt15.
+This comparison confirms expm_flow_opt15 is the fastest implementation, with expm_flow exhibiting the largest runtime overhead.
 Notably, these execution time ratios are lower than the corresponding matrix product ratios.
@@
 In this subsection, the results of tests performed directly on the code available in [23] will be described.
-As mention before, this code incorporates a matrix exponential function into neural networks implementing generative flows models.
+As mention before, this code incorporates a matrix exponential function into neural networks implementing generative flow models.
 In our executions, the matrix exponential is computed using the following functions implemented in PyTorch:
   1. 1.
-    expm_flow_ps is based on Algorithm 2 and evaluates the polynomial using the Paterson–Stockmeyer method.
+    expm_flow_ps is based on Algorithms 2 and $3.$ The polynomial is evaluated using the Paterson–Stockmeyer method.
   2. 2.
-    expm_flow_sastre: Analogous to the previous function, this implementation is based on Algorithm 2, but the polynomials are evaluated using formulas (10)-(17).
+    expm_flow_opt15: Analogous to the previous function, this implementation is based on Algorithms 2 and 3.3, where the polynomials are evaluated using formulas (10)-(17).
   3. 3.
     expm_flow directly corresponds to the matrix exponential approximation method presented in [23].
-All executions were carried out using CUDA and the graphics card of the machine involved.
-As testbeds in training, CIFAR-10 [37], ImageNet32 and ImageNet64 [38] datasets were used.
-The number of invocations analysed to extract results to the functions responsible for calculating the matrix exponential during the training process was 5000.
-For each call to the exponential function, multiple data were recorded for the matrices comprising the tensor to be computed, such as the number of matrices, their sizes, and the highest matrix $\infty$-norm of all of them.
-Thus, we can confirm that the $\infty$-norm of the matrices studied varied from 2.84e-4 to 12.57 for CIFAR-10, from 1.17e-05 to 12.49 for ImageNet32, and from 1.27e-05 to 12.8 for ImageNet64.
+To maintain consistency with the baseline implementation expm_flow [23], both expm_flow_ps and expm_flow_opt15 utilize the $\infty$-norm.
+All computations were performed using NVIDIA CUDA on a dedicated GPU.
+The proposed method was evaluated during the training process using the CIFAR-10 [35], ImageNet32, and ImageNet64 [36] datasets.
+To obtain representative performance metrics, we analyzed 5000 invocations of the matrix exponential function.
+For each invocation, we recorded metadata for the matrices comprising the input tensor, specifically the number of matrices in the batch, their dimensions, and the maximum $\infty$-norm across the set.
+The $\infty$-norms of the matrices studied range from $2.84\times 10^{-4}$ to $12.57$ for CIFAR-10, from $1.17\times 10^{-5}$ to $12.49$ for ImageNet32, and from $1.27\times 10^{-5}$ to $12.8$ for ImageNet64.
+For a normal matrix with the minimum observed norm of $1.27\times 10^{-5}$, the baseline Algorithm 1 requires a Taylor polynomial of degree $m=7$ with a cost of $6M$ matrix products according to (7).
+In contrast, Algorithm 3.3 utilizes a degree $m=8$ with a cost of only $3M$.
+Since the scaling parameter is $s=0$ for both, the baseline method theoretically requires twice as many matrix products as the proposed approach.
+Similarly, for a normal matrix with the maximum observed norm of $12.8$, the baseline Algorithm 1 selects $m=8$ and $s=5$, leading to a total cost of $12M$ matrix products.
+In contrast, Algorithm 3.3 utilizes $m=15$ with $s=3$, which reduces the cost to $7M$.
+This implies that the baseline method is theoretically $1.71$ times more expensive than the proposed approach for this case.
 In addition, the following metrics were logged for each of the three comparison functions: the degree of the polynomial used, the scaling parameter, the maximum error observed across all matrices in the tensor, the total number of matrix products involved, and the required response time.
@@
 Figures 2(a), 3(a), and 4(a) plot the normwise relative errors just for the first 150 training runs.
-For clarity, these same results are shown sorted, but now for the 5000 runs, from largest to smallest error independently for the three functions analysed in Figures 2(b), 3(b), and 4(b).
+For clarity, these same results are shown sorted, but now for the 5000 runs, from largest to smallest error independently for the three functions analyzed in Figures 2(b), 3(b), and 4(b).
 Unlike 1(b), there is here no correspondence between a matrix on the $x$-axis and the three errors that appear on the $y$-axis, due to the aforementioned sorting.
 As we can observe, expm_flow always resulted in higher errors in the exponential approximation.
-In contrast, expm_flow_sastre and expm_flow_ps provided very similar results, with expm_flow_sastre being slightly more accurate for the three datasets.
-This conclusion can also be deduced from the Figures 2(c), 3(c), and 4(c) related to the performance profile, where expm_flow_sastre and expm_flow_ps quickly provided probabilities equal to 1.
-Figures2(d), 3(d), and 4(d) also confirm that code expm_flow_sastre delivered a more accurate result than the other functions in a higher percentage of cases, with values of 50%, 49%, and 50% of the matrices for the three datasets, respectively.
-In turn, function expm_flow_sastre provided similar but slightly lower percentages, with values of 48%, for CIFAR-10, and 45%, for ImageNet32 and ImageNet64.
+In contrast, expm_flow_opt15 and expm_flow_ps provided very similar results, with expm_flow_opt15 being slightly more accurate for the three datasets.
+This conclusion can also be deduced from the Figures 2(c), 3(c) and 4(c) related to the performance profile, where expm_flow_opt15 and expm_flow_ps quickly provided probabilities equal to 1.
+Figures2(d), 3(d), and 4(d) also confirm that code expm_flow_opt15 delivered a more accurate result than the other functions in a higher percentage of cases, with values of 50%, 49%, and 50% of the matrices for the three datasets, respectively.
+In turn, function expm_flow_opt15 provided similar but slightly lower percentages, with values of 48%, for CIFAR-10, and 45%, for ImageNet32 and ImageNet64.
 In clear accordance with Figures 1(e), Figures 2(e), 3(e), and 4(e) also reveal that expm_flow used polynomials of a lower order than those of the other functions.
 In contrast, expm_flow required higher values of the scaling parameter, as can be appreciated in Figures 2(f), 3(f), and 4(f).
-Finally, with regard to the number of products, Figures 2(g), 3(g), and 4(g) state that expm_flow carried out 1.99, 1.86, and 1.88 times more products than expm_flow_sastre in the calculation of matrix exponentials.
-Similarly, Figures 2(h), 3(h), and 4(h) show that the exponential computation based on function expm_flow_sastre was 1.87, 1.97, and 2.5 times faster than the same calculation using function expm_flow.
+Finally, with regard to the number of products, Figures 2(g), 3(g), and 4(g) state that expm_flow carried out 1.99, 1.86, and 1.88 times more products than expm_flow_opt15 in the calculation of matrix exponentials.
+Similarly, Figures 2(h), 3(h), and 4(h) show that the exponential computation based on function expm_flow_opt15 was 1.87, 1.97, and 2.5 times faster than the same calculation using function expm_flow.
 (a) Normwise relative error.
@@
 5 Performance analysis in generative flow models
-To evaluate the efficiency gains of the proposed matrix exponential approach (expm_flow_sastre) relative to the baseline (expm_flow), we conducted a series of numerical experiments using three standard computer vision datasets: CIFAR-10, ImageNet32, and ImageNet64.
+To evaluate the efficiency gains of the proposed matrix exponential approach (expm_flow_opt15) relative to the baseline (expm_flow), we conducted a series of numerical experiments on the three aforementioned datasets: CIFAR-10, ImageNet32, and ImageNet64.
 All tests were executed on a system equipped with an Intel Xeon Platinum 8268 Processor, an NVIDIA RTX8000 GPU and 128 GB of RAM.
+Note that in this section both the matrix exponential and the reduced-rank approximation (8) are used following the generative model framework in [23].
+Specifically, the matrix exponential is utilized for matrices smaller than $24\times 24$, while the approximation (8) is applied to larger matrices.
+In the context of the reduced-rank approximation, where no scaling is applied ($s=0$), the computational cost of the baseline algorithm increases linearly with the degree $m$ according to (7).
+Consequently, the proposed algorithm is significantly more efficient, as demonstrated in Section 3.1 for Taylor based evaluation formula with approximation order $m=15+$.
+Furthermore, this performance gap widens as the polynomial degree increases.
+Moreover, as suggested by Remark 1, these efficiency ratios could be more significant for nonnormal matrices.
 The evaluation consists of two parts.
@@
 Second, both algorithms were integrated into the coupling layers of a Glow-based architecture [37], following the methodology described in [23].
-Performance is quantified by measuring total computation time and calculating the speed-up ratio, defined as the execution time of expm_flow divided by that of expm_flow_sastre.
+Performance is quantified by measuring total computation time and calculating the speed-up ratio, defined as the execution time of expm_flow divided by that of expm_flow_opt15.
 The flow model architecture, illustrated in Figure 5, follows a multi-scale structure.
@@
 5.1 Computational scaling and matrix dimensionality
-This analysis evaluates how the performance of the proposed matrix exponential implementation, expm_flow_sastre, scales relative to the baseline, expm_flow, as matrix dimensionality increases.
+This analysis evaluates how the performance of the proposed matrix exponential implementation, expm_flow_opt15, scales relative to the baseline, expm_flow, as matrix dimensionality increases.
 We examine matrix orders $n\in\{8,16,32,64,128,256,512,1024\}$ under two distinct scenarios: individual matrix transformations of size $n\times n$ and batched tensor operations of size $n\times 16\times 16$.
@@
 Consequently, the proposed method yields substantial benefits by reducing the required multiplication count, effectively leveraging the parallel linear algebra routines provided by the PyTorch backend, such as cuBLAS and MKL.
-Figure 6: Comparison of the execution time required to calculate the matrix exponential using the original method (expm_flow) and the proposed method (expm_flow_sastre).
+Figure 6: Comparison of the execution time required to calculate the matrix exponential using the original method (expm_flow) and the proposed method (expm_flow_opt15).
 The results are presented for different matrix sizes $n\in\{2,4,8,16,32,64,128,256,512,1024\}$.
@@
 Glow [37] is a coupling layer approach to learning invertible transformations.
-We replaced the original implementation (expm_flow) with our optimized function (expm_flow_sastre) to measure its impact on computational efficiency during the training phase.
-All models were trained for 50 epochs across three benchmark datasets: CIFAR-10, ImageNet32, and ImageNet64.
+We replaced the original implementation (expm_flow) with our optimized function (expm_flow_opt15) to measure its impact on computational efficiency during the training phase.
+All models were trained for 50 epochs across the three aforementioned benchmark datasets CIFAR-10, ImageNet32, and ImageNet64.
 As detailed in Table 6, the proposed method demonstrates increased computational efficiency across all tested configurations.
 While the minimum observed speed-up was $3.91$ for ImageNet64, the most significant gain occurred with ImageNet32, where training was nearly 10 times faster ($9.74$) than the baseline.
-This dramatic reduction in training time, specifically the shift from 32,304 seconds to 3,317 seconds per epoch on ImageNet32, demonstrates that the computational bottleneck previously associated with matrix exponentials in flow-based models can be effectively mitigated.
+This dramatic reduction in training time, specifically the shift from 32304 seconds to 3317 seconds per epoch on ImageNet32, shows that the computational bottleneck previously associated with matrix exponentials in flow-based models can be effectively mitigated.
 Dataset CIFAR-10 ImageNet32 ImageNet64
 expm_flow time 2463 32304 42482
-expm_flow_sastre time 443.4 3317 10875
+expm_flow_opt15 time 443.4 3317 10875
 Speed-up 5.554 9.739 3.906
 Table 6: Average training time per epoch (in seconds) for the matrix exponential version of Glow.
-Results compare the efficiency of our proposed expm_flow_sastre function against the original expm_flow implementation across three standard datasets.
+Results compare the efficiency of our proposed expm_flow_opt15 function against the original expm_flow implementation across three standard datasets.
 5.3 Inference and Sampling Performance
 While the previous sections evaluated performance during the training phase, efficiency during inference is equally critical for generative models.
-Table 5 presents a comparison of the sampling times between expm_flow and expm_flow_sastre.
+Table 7 presents a comparison of the sampling times between expm_flow and expm_flow_opt15.
 In this context, sampling time refers to the duration required for the model to generate an image from the latent space.
@@
 The results demonstrate that the speedup provided by our method becomes significantly more pronounced as the batch size increases.
-While the performance is nearly identical for a single sample, the expm_flow_sastre variant achieves a speedup factor of approximately $1.95$ for larger batches, effectively halving the inference latency.
+While the performance is nearly identical for a single sample, the expm_flow_opt15 variant achieves a speedup factor of approximately $1.95$ for larger batches, effectively halving the inference latency.
 Sample 1 sample 128 samples
 expm_flow time 0.3839 0.8286
-expm_flow_sastre time 0.3836 0.4247
+expm_flow_opt15 time 0.3836 0.4247
 Speed-up 1.001 1.951
 Table 7: Average inference times (in seconds) for sampling a matrix exponential variant of Glow.
-We compare the original expm_flow against expm_flow_sastre for single and batched samples, including the calculated speedup factor.
+We compare the original expm_flow against expm_flow_opt15 for single and batched samples, including the calculated speedup factor.
 6 Conclusions
@@
 Numerical evaluations conducted on both a diverse testbed of matrices and practical generative models demonstrate that, relative to the original implementation, the proposed Taylor-based approach yields training speedups between $3.91$ and $9.74$ while reducing inference latency by approximately 50% in large-batch scenarios.
-These results demonstrate that modern polynomial evaluation strategies allow for higher-order approximations at a lower computational cost, measured in terms of matrix multiplications ($M$), than traditional rational or Paterson–Stockmeyer-based methods.
-Consequently, the proposed implementation offers a portable, library-independent solution for high-throughput applications and large-scale simulation environments.
+These results demonstrate that modern polynomial evaluation strategies allow for higher-order approximations at a lower computational cost, measured in terms of matrix multiplications ($M$), than traditional Paterson–Stockmeyer-based methods.
+Consequently, the proposed implementation offers a portable, library-independent solution for high-throughput applications.
 Acknowledgements